### v1

# Reduced Network Cumulative Constraint Violation for Distributed Bandit Convex Optimization under Slater’s Condition

Abstract
This paper studies the distributed bandit convex optimization problem with time-varying inequality constraints, where the goal is to minimize network regret and cumulative constraint violation. To calculate network cumulative constraint violation, existing distributed bandit online algorithms solving this problem directly use the clipped constraint function to replace its original constraint function. However, the use of the clipping operation renders Slater’s condition (i.e, there exists a point that strictly satisfies the inequality constraints at all iterations) ineffective to achieve reduced network cumulative constraint violation. To tackle this challenge, we propose a new distributed bandit online primal–dual algorithm. If local loss functions are convex, we show that the proposed algorithm establishes an $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ network regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ network cumulative constraint violation bound, where $T$ is the total number of iterations and $c\in({0,1})$ is a user-defined trade-off parameter. When Slater’s condition holds, the network cumulative constraint violation bound is reduced to $\mathcal{O}({{T^{1-c}}})$. In addition, if local loss functions are strongly convex, for the case where strongly convex parameters are unknown, the network regret bound is reduced to $\mathcal{O}({{T^{1-c}}})$, and the network cumulative constraint violation bound is $\mathcal{O}({{T^{1-c/2}}})$ and $\mathcal{O}({{T^{1-c}}})$ without and with Slater’s condition, respectively. For the case where strongly convex parameters are known, the network regret bound is further reduced to $\mathcal{O}\big{(}{\log(T)}\big{)}$, and the network cumulative constraint violation bound is reduced to $\mathcal{O}\big{(}{\sqrt{\log(T)T}}\big{)}$ and $\mathcal{O}\big{(}{\log(T)}\big{)}$ without and with Slater’s condition, respectively. To the best of our knowledge, this paper is among the first to establish reduced (network) cumulative constraint violation bounds for (distributed) bandit convex optimization with time-varying constraints under Slater’s condition. Finally, a numerical example is provided to verify the theoretical results.
Index Terms:
Bandit convex optimization, cumulative constraint violation, distributed optimization, Slater’s condition, time-varying constraints.
I INTRODUCTION
Bandit convex optimization is a sequential decision process in dynamic environments, which can be understood as a structured repeated game with $T$ iterations between a decision maker and an adversary [1]. Specifically, at iteration $t$, a decision maker chooses $x_{t}$ from a convex set $\mathbb{X}\in{\mathbb{R}^{p}}$ in an Euclidean space. After committing to this choice, she receives one-point bandit feedback for a convex loss function ${f_{t}}:\mathbb{X}\to\mathbb{R}$ from the adversary (i.e., the value of the function ${f_{t}}$ at $x_{t}$ is revealed to the decision maker by the adversary), where $\mathbb{R}$ denotes the set of all real numbers. Accordingly, the decision maker suffers a loss ${f_{t}}({{x_{t}}})$. The goal of the decision maker is to minimize the accumulative loss $\sum\nolimits_{t=1}^{T}{{f_{t}}({{x_{t}}})}$ over $T$ iterations. The standard measure metric is the regret
$\sum\limits_{t=1}^{T}{{f_{t}}({{x_{t}}})}-\mathop{\min}\limits_{{}_{x\in\mathbb{X}}}\sum\limits_{t=1}^{T}{{f_{t}}(x)},$
which measures the difference between the accumulative loss and the loss induced by the best fixed decision in hindsight. Over the past decades, bandit convex optimization has garnered substantial interest, see, e.g., [2] and references therein, due to its wide applications including smart grids with uncertain supply of renewable energy [3, 4] and data centers with uncertain user demands [5, 6, 7].
Various bandit online algorithms with sublinear regret have been developed. For example, in [8], the authors propose a bandit online projection gradient descent where the gradient of the loss function is approximated by using a one-point estimate, establishing an ${\cal O}({T^{3/4}})$ regret bound for general convex loss functions. The bound is further reduced to $\mathcal{O}({{T^{2/3}}})$ for more stringent strong convex loss functions in [9]. However, a lower bound from [10] implies that the regret of this algorithm in [8] will be $\Omega({\sqrt{T}})$, even for strongly convex loss functions. The lower bound is much worse than the $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound established by the algorithm in [11] for the full-information feedback setting (the convex loss function is revealed to the decision maker at each iteration). To deal with this challenge, the authors of [12] extend bandit convex optimization by allowing that the values of the convex loss function at multiple points are simultaneously revealed to the decision maker. Moreover, the authors propose a bandit online projection gradient descent where the gradient of the loss function is approximated by using two-point estimate, and establish an $\tilde{\mathcal{O}}({\sqrt{T}})$ regret bound and an $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound for general convex loss functions and strong convex loss functions, respectively. These bounds closely resemble the optimal bounds (i.e., the $\mathcal{O}({\sqrt{T}})$ regret bound established by the algorithm in [13] for general convex loss functions and the $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound established by the algorithm in [11] for strong convex loss functions) for the full-information feedback setting.
Note that the aforementioned algorithms require the projection operator onto the feasible set at each iteration. The operator would be straightforward if the feasible set is a simple set, e.g., a cube, a ball, or a simplex, while it would yield heavy computation burden if the feasible set is complicated. In practice, the feasible set is often characterized by inequality constraints, i.e.,
$\mathcal{X}=\{{x:g(x)\leq{\mathbf{0}_{m}},x\in\mathbb{X}}\},$
where $g(x):{\mathbb{R}^{p}}\to{\mathbb{R}^{m}}$ is the static convex constraint function, $m$ is a positive integer, and $\mathbb{X}$ is normally a simple set. In this case, long term constraints are considered in [14] for bandit convex optimization, where decisions are chosen from the simple set $\mathbb{X}$ and static inequality constraints should be satisfied in the long term on average. To measure accumulative violation of inequality constraints, a performance metric, constraint violation, is defined as
$\Big{\|}{{{\Big{[}{\sum\limits_{t=1}^{T}{g({{x_{t}}})}}\Big{]}_{+}}}}\Big{\|},$ (1)
[…]
Distributed paradigm presents a promising framework for overcoming the limitations of centralized ones including single point of the failure, data privacy and heavy computation overhead [16, 17]. Therefore, distributed bandit convex optimization is increasingly garnering significant attention, see [18, 19, 20, 21, 22, 23, 24]. In this problem, the loss function ${f_{t}}(x)$ at each iteration $t$ is decomposed across a network of $n$ agents by ${f_{t}}(x)=\frac{1}{n}\sum\nolimits_{i=1}^{n}{{f_{i,t}}(x)}$, where ${{f_{i,t}}}$ is called the local loss function. Each agent chooses its own decision ${x_{i,t}}$ from the set $\mathbb{X}$, and then the values of its local loss function ${{f_{i,t}}}$ at some points are revealed to itself only by the adversary. The goal of the agents is to minimize the network-wide accumulated loss, and the corresponding performance metric can be the network regret
$\frac{1}{n}\sum\limits_{i=1}^{n}{\Big{(}{\sum\limits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}-\mathop{\min}\limits_{x\in X}\sum\limits_{t=1}^{T}{{f_{t}}({{x}})}}\Big{)}}.$
Recently, the authors of [25] extend this problem by using the idea of long term constraints, and use a new form of constraint violation metric proposed in [26]. This metric is called cumulative constraint violation, and is given by
$\Big{\|}{\sum\limits_{t=1}^{T}{{{[{g({{x_{t}}})}]}_{+}}}}\Big{\|}.$ (2)
The constraint violation defined in (1) takes the summation across iterations before the projection operation ${[\cdot]_{+}}$ such that it allows strict feasible decisions having large margins compensate constraint violations at many iterations, cumulative constraint violation defined in (2) considers all constraints that are not satisfied, and thus cumulative constraint violation is stricter than constraint violation. To calculate the network-wide accumulated cumulative constraint violation, the corresponding performance metric can be the network cumulative constraint violation
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g}({{x_{i,t}}})}]}_{+}}}\|}}.$
When local loss functions are quadratic and constraint functions are linear, the authors of [25] establish an $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ network regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ network cumulative constraint violation bound with $c\in({0,1})$. For more general convex local loss and constraint functions, the authors of [27] propose a distributed online primal–dual algorithm with one-point bandit feedback, and establish an $\mathcal{O}\big{(}{{T^{\max\{{c,1-c/3}\}}}}\big{)}$ network regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ network cumulative constraint violation bound with $c\in({0,1})$. When local loss functions are strongly convex, an $\mathcal{O}\big{(}{{T^{2/3}}\log(T)}\big{)}$ network regret bound and an $\mathcal{O}\big{(}{\sqrt{T\log(T)}}\big{)}$ network cumulative constraint violation bound are established. The authors of [28] further extend this problem to the time-varying constraints setting, where the time-varying convex constraint function is denoted by $g_{t}(x):{\mathbb{R}^{p}}\to{\mathbb{R}^{m}}$. Moreover, they propose a distributed online primal–dual algorithm with two-point bandit feedback, and establish a reduced $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ network regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ network cumulative constraint violation bound for general convex loss functions, and a reduced $\mathcal{O}({{T^{c}}})$ network regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ for strong convex loss functions, where $c\in({0,1})$. For calculating (network) cumulative constraint violation, a key idea in [25, 26, 27, 28] is to use the clipped constraint function ${[g]_{+}}$ or ${[{{g_{t}}}]_{+}}$ to replace the original constraint function $g$ or ${{g_{t}}}$, respectively. However, in this way, reduced network cumulative constraint violation bounds cannot be established under Slater’s condition [29], i.e., the use of the clipping operation renders Slater’s condition ineffective. Slater’s condition is a sufficient condition for strong duality to hold in convex optimization problems [30], which is used to establish reduced constraint violation for the full-information feedback setting in [31, 32]. Moreover, the authors of [29] propose a distributed online primal–dual algorithm with full-information feedback where the clipped constraint function ${[{{g_{t}}}]_{+}}$ is not directly used to replace the original constraint function ${{g_{t}}}$. Instead, the algorithm updates the dual variables by directly maximizing the regularized Lagrangian function. In particular, with this idea, Slater’s condition is still effective.
Note that such a distributed bandit online algorithm that can achieve reduced network cumulative constraint violation under Slater’s condition is still missing, which motivates our study. In particular, in this paper, we consider the distributed bandit convex optimization problem with time-varying constraints where the decision makers receive bandit feedback for both loss and constraint functions at each iteration, and use network regret and cumulative constraint violation as performance metrics. The contributions are summarized as follows.
  • $\bullet$
    This paper proposes a new distributed bandit online primal–dual algorithm. Different from the distributed bandit online algorithm in [28] where the clipped constraint function ${[{{g_{t}}}]_{+}}$ is directly used to replace the original constraint function ${{g_{t}}}$, the proposed algorithm updates the dual variables by directly maximizing the regularized Lagrangian function, which can be explicitly calculated using the clipped constraint function. Different from the distributed online algorithm in [29] where the dual variables are updated by using the composite objective mirror descent and the subgradients of local loss and constraint functions are directly used, the proposed algorithm updates the dual variables by using the projected gradient descent instead of the composite objective mirror descent used in [29]. Moreover, the proposed algorithm uses two-point stochastic estimators to approximate these subgradients as they are unavailable in the bandit feedback setting. Note that the gaps between the estimators and the subgradients cause nontrivial challenges for performance analysis, which will be explained in detail in Remark 2. More importantly, the proposed algorithm enables the use of the stricter cumulative constraint violation metric while preserving the effectiveness of Slater’s condition.
  • $\bullet$
    For convex local loss functions, we show in Theorem 1 that the proposed algorithm establishes an $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ network regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ network cumulative constraint violation bound with $c\in({0,1})$, which are the same as the results established in [28, 29], generalize the results established in [14, 15, 25], and also improve the results established in [27]. When Slater’s condition holds, we show in Theorem 2 that the network cumulative constraint violation bound is reduced to $\mathcal{O}({{T^{1-c}}})$. To the best of our knowledge, this paper is among the first to establish a reduced cumulative constraint violation bound for bandit convex optimization with long term constraints under Slater’s condition.
  • $\bullet$
    For strongly convex local loss functions, under unknown strongly convex parameters, we show in Theorem 3 that the proposed algorithm establishes an $\mathcal{O}({{T^{1-c}}})$ network regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ network cumulative constraint violation bound with $c\in({0,1})$, moreover, the network cumulative constraint violation bound is reduced to $\mathcal{O}({{T^{1-c}}})$ when Slater’s condition holds. Under known strongly convex parameters, we show in Theorem 4 that the proposed algorithm establishes an $\mathcal{O}\big{(}{\log(T)}\big{)}$ network regret bound and an $\mathcal{O}\big{(}{\sqrt{\log(T)T}}\big{)}$ network cumulative constraint violation bound, which are the same as the results established in [29], and improve the results established in [27, 28]. Moreover, the network cumulative constraint violation bound is reduced to $\mathcal{O}\big{(}{\log(T)}\big{)}$ when Slater’s condition holds. Note that this paper is among the first to establish such a result for (distributed) bandit convex optimization with long term constraints.
TABLE I: Comparison of this paper to related works on bandit convex optimization
with long term constraints.
Reference Problem type Loss and constraint functions Slater’s condition Information feedback Regret Constraint violation Cumulative constraint violation
[14] Centralized Convex, convex and static $\times$ $\nabla{f_{t}}$, and two-point bandit feedback for ${g}$ $\mathcal{O}({\sqrt{T}})$ $\mathcal{O}({{T^{3/4}}})$ Not given
[15] Centralized Convex, convex and time-varying $\times$ Two-point bandit feedback for ${f_{t}}$ and ${g_{t}}$ $\mathcal{O}({\sqrt{T}})$ $\mathcal{O}({{T^{3/4}}})$ Not given
[25] Distributed Quadratic, linear and static $\times$ $\nabla{g}$, and two-point bandit feedback for ${f_{t}}$ $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ $\mathcal{O}({{T^{1-c/2}}})$
[27] Distributed Convex, convex and static $\times$ $\nabla{g}$, and one-point bandit feedback for ${f_{t}}$ $\mathcal{O}\big{(}{{T^{\max\{{c,1-c/3}\}}}}\big{)}$ $\mathcal{O}({{T^{1-c/2}}})$
Strongly convex, convex and static $\mathcal{O}\big{(}{{T^{2/3}}\log(T)}\big{)}$ $\mathcal{O}\big{(}{\sqrt{T\log(T)}}\big{)}$
[28] Distributed Convex, convex and time-varying $\times$ Two-point bandit feedback for ${f_{t}}$ and ${g_{t}}$ $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ $\mathcal{O}({{T^{1-c/2}}})$
Strongly convex, convex and time-varying $\mathcal{O}({{T^{c}}})$ $\mathcal{O}({{T^{1-c/2}}})$
This paper Distributed Convex, convex and time-varying $\times$ Two-point bandit feedback for ${f_{t}}$ and ${g_{t}}$ $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ $\mathcal{O}({{T^{1-c/2}}})$
$\checkmark$ $\mathcal{O}({{T^{1-c}}})$
Strongly convex, convex and time-varying $\times$ $\mathcal{O}\big{(}{\log(T)}\big{)}$ $\mathcal{O}\big{(}{\sqrt{\log(T)T}}\big{)}$
$\checkmark$ $\mathcal{O}\big{(}{\log(T)}\big{)}$
The detailed comparison of this paper to related studies is summarized in TABLE I, where we only present the static part of regret for the sake of clarity.
The remainder of this paper is organised as follows. Section II presents the problem formulation. Section III proposes the distributed bandit online primal–dual algorithm, and analyzes its performance. Section IV demonstrates a numerical simulation to verify the theoretical results. Finally, Section V concludes this paper. All proofs are given in Appendix.
Notations: ${\mathbb{N}_{+}}$, $\mathbb{R}$, ${\mathbb{R}^{p}}$ and $\mathbb{R}_{+}^{p}$ denote the sets of all positive integers, real numbers, $p$-dimensional and nonnegative vectors, respectively. Given $m$ and $n\in{\mathbb{N}_{+}}$, $[m]$ denotes the set $\{{1,\cdot\cdot\cdot,m}\}$, and $[m,n]$ denotes the set $\{{m,\cdot\cdot\cdot,n}\}$ for $m<n$. Given vectors $x$ and $y$, ${x^{T}}$ denotes the transpose of the vector $x$, and $\langle{x,y}\rangle$ and $x\otimes y$ denote the standard inner and Kronecker product of the vectors $x$ and $y$, respectively. ${\mathbf{0}_{m}}$ denotes the $m$-dimensional column vector whose components are all 00. $\mathrm{col}({q_{1}},\cdot\cdot\cdot,{q_{n}})$ denotes the concatenated column vector of ${q_{i}}\in{\mathbb{R}^{{m_{i}}}}$ for $i\in[n]$. ${\mathbb{B}^{p}}$ and ${\mathbb{S}^{p}}$ denote the unit ball and sphere centered around the origin in ${\mathbb{R}^{p}}$ under Euclidean norm, respectively. $\mathbf{E}$ denotes the expectation. For a set $\mathbb{K}\in{\mathbb{R}^{p}}$ and a vector $x\in{\mathbb{R}^{p}}$, ${\mathcal{P}_{\mathbb{K}}}(x)$ denotes the projection of the vector $x$ onto the set $\mathbb{K}$, i.e., ${\mathcal{P}_{\mathbb{K}}}(x)=\arg{\min_{y\in{\mathbb{K}}}}{\|{x-y}\|^{2}}$, and $[x]_{+}$ denotes ${\mathcal{P}_{\mathbb{R}_{+}^{p}}}(x)$. For a function $f$ and a vector $x$, $\partial f(x)$ denotes the subgradient of $f$ at $x$.
II PROBLEM FORMULATION
Consider the distributed bandit convex optimization problem with time-varying constraints. At iteration $t$, a network of $n$ agents is modeled by a time-varying directed graph ${\mathcal{G}_{t}}=({\mathcal{V},{\mathcal{E}_{t}}})$ with the agent set $\mathcal{V}=[n]$ and the edge set ${\mathcal{E}_{t}}\subseteq\mathcal{V}\times\mathcal{V}$. $({j,i})\in{\mathcal{E}_{t}}$ indicates that agent $i$ can receive information from agent $j$. The sets of in- and out-neighbors of agent $i$ are $\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})=\{{j\in[n]|({j,i})\in{\mathcal{E}_{t}}}\}$ and $\mathcal{N}_{i}^{\text{out}}({{\mathcal{G}_{t}}})=\{{j\in[n]|({i,j})\in{\mathcal{E}_{t}}}\}$, respectively. An adversary first erratically selects $n$ convex functions $\{{{f_{i,t}}:\mathbb{X}\to\mathbb{R}}\}$ and $n$ convex constraint functions $\{{{g_{i,t}}:\mathbb{X}\to{\mathbb{R}^{{m_{i}}}}}\}$ for $i\in[n]$, where $\mathbb{X}\subseteq{\mathbb{R}^{p}}$ is a known convex set, and both ${m_{i}}$ and $p$ are positive integers. Then, the agents collaborate to select their local decisions $\{{{x_{i,t}}\in\mathbb{X}}\}$ without prior access to $\{{{f_{i,t}}}\}$ and $\{{{g_{i,t}}}\}$. At the same time, the values of ${f_{i,t}}$ and ${g_{i,t}}$ at the point ${x_{i,t}}$ as well as at other potential points are privately revealed to each agent $i$. The goal of the agents is to choose the decision sequence $\{{{x_{i,t}}}\}$ for $i\in[n]$ and $t\in[T]$ such that both network regret
${\rm{Net}\mbox{-}\rm{Reg}}(T)$ $:=\frac{1}{n}\sum\limits_{i=1}^{n}{\Big{(}{\sum\limits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}-\mathop{\min}\limits_{x\in{\mathcal{X}_{T}}}\sum\limits_{t=1}^{T}{{f_{t}}(x)}}\Big{)}},$ (3)
and network cumulative constraint violation
${\rm{Net}\mbox{-}\rm{CCV}}(T)$ $:=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}},$ (4)
increase sublinearly, where ${f_{t}}(x)=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{f_{j,t}}(x)}$ is the global loss function of the network at iteration $t$, ${g_{t}}(x)={\rm{col}}\big{(}{{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)}\big{)}\in{\mathbb{R}^{m}}$ with $m=\sum\nolimits_{i=1}^{n}{{m_{i}}}$ is the global constraint function of the network at iteration $t$, and
${\mathcal{X}_{T}}=\{{x:x\in\mathbb{X},{g_{t}}(x)\leq{\mathbf{0}_{m}},\forall t\in[T]}\},$ (5)
is the feasible set. To guarantee that the offline optimal static decision always exists, we assume that for any $T\in{\mathbb{N}_{+}}$, the feasible set ${\mathcal{X}_{T}}$ is nonempty.
Note that when ${g_{i,t}}\equiv{\mathbf{0}_{{m_{i}}}}$, $\forall i\in[n]$, $\forall t\in{\mathbb{N}_{+}}$, the considered distributed problem becomes the problem studied in [19, 20, 21, 22, 23, 24]; when ${g_{i,t}}\equiv g$, $\forall i\in[n]$, $\forall t\in{\mathbb{N}_{+}}$ with $g$ being a known constraint function, the considered distributed problem becomes the problem studied in [25, 27].
The following commonly used assumption on the set $\mathbb{X}$ is made.
Assumption 1.
The set $\mathbb{X}$ is closed. Moreover, the convex set $\mathbb{X}$ contains the ball of radius $r(\mathbb{X})$ and is contained in the ball of radius $R(\mathbb{X})$, i.e.,
$r(\mathbb{X}){\mathbb{B}^{p}}\subseteq\mathbb{X}\subseteq R(\mathbb{X}){\mathbb{B}^{p}}.$ (6)
The following assumptions on the loss and constraint functions are made.
Assumption 2.
For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, there exists a constant ${F}$ such that
$|{{f_{i,t}}(x)}|$ $\leq{F}.$ (7)
Assumption 3.
For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, the functions ${f_{i,t}}$ and ${{g_{i,t}}}$ are convex, and the subgradients $\partial{f_{i,t}}(x)$ and $\partial{g_{i,t}}(x)$ exist. Moreover, there exist constants ${G_{1}}$ and ${G_{2}}$ such that
$\|{\partial{f_{i,t}}(x)}\|$ $\leq{G_{1}},$ (8a)
$\|{\partial{g_{i,t}}(x)}\|$ $\leq{G_{2}},x\in\mathbb{X}.$ (8b)
Note that we do not need the assumption that the local constraint functions $\{{{g_{i,t}}}\}$ are uniformly bounded while [28] needs it. In addition, from Assumption 3, and Lemma 2.6 in [33], for all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, we have
$|{{f_{i,t}}(x)-{f_{i,t}}(y)}|$ $\leq{G_{1}}\|{x-y}\|,$ (9a)
$\|{{g_{i,t}}(x)-{g_{i,t}}(y)}\|$ $\leq{G_{2}}\|{x-y}\|,x,y\in\mathbb{X}.$ (9b)
The following assumption on the graph is made, which is also used in [34, 28, 29, 35].
Assumption 4.
For $t\in{\mathbb{N}_{+}}$, the time-varying directed graph $\mathcal{G}_{t}$ satisfies that
(i) There exists a constant $w\in({0,1})$ such that ${[{{W_{t}}}]_{ij}}\geq w$ if $({j,i})\in{\mathcal{E}_{t}}$ or $i=j$, and ${[{{W_{t}}}]_{ij}}=0$ otherwise.
(ii) The mixing matrix ${W_{t}}$ is doubly stochastic, i.e., ${\sum\nolimits_{i=1}^{n}{[{{W_{t}}}]}_{ij}}={\sum\nolimits_{j=1}^{n}{[{{W_{t}}}]}_{ij}}=1$, $\forall i,j\in[n]$.
(iii) There exists an integer $B>0$ such that the time-varying directed graph $({\mathcal{V},\cup_{l=0}^{B-1}{\mathcal{E}_{t+l}}})$ is strongly connected.
As stated in the introduction, compared to [28], this paper establishes reduced network cumulative constraint violation bounds under Slater’s condition. In the following, we formally introduce Slater’s condition.
Assumption 5.
(Slater’s condition) There exists a point ${x_{s}}\in\mathbb{X}$ and a positive constant ${\varsigma_{s}}$ such that
${g_{t}}({{x_{s}}})\leq-{\varsigma_{s}}{\mathbf{1}_{m}},t\in{\mathbb{N}_{+}}.$ (10)
Slater’s condition is a sufficient condition for strong duality to hold in convex optimization problems [30]. To the best of our knowledge, there are no studies to show that reduced cumulative constraint violation bounds can be established in bandit convex optimization. To calculate network cumulative constraint violation, [28] directly replaces the original constraint functions with the corresponding clipped constraint functions, which makes Slater’s condition ineffective. In this paper, we propose a new distributed bandit online algorithm where Slater’s condition remains effective.
III DISTRIBUTED BANDIT ONLINE PRIMAL–DUAL ALGORITHM
III-A Algorithm description
For the global loss function ${f_{t}}$ and constraint function ${g_{t}}$, the associated regularized Lagrangian function is
${\mathcal{L}_{t}}({{x_{t}},{q_{t}}}):=\frac{1}{n}\sum\limits_{i=1}^{n}{{f_{i,t}}({{x_{t}}})}+q_{t}^{T}{g_{t}}({{x_{t}}})-\frac{1}{{2{\gamma_{t}}}}{\|{{q_{t}}}\|^{2}},$ (11)
where ${x_{t}}\in{\mathbb{R}^{p}}$ and ${q_{t}}\in\mathbb{R}_{+}^{m}$ represent the primal and dual variables, respectively, and ${{\gamma_{t}}}$ is the regularization parameter. The primal and dual variables can be updated by the standard projected primal–dual algorithm
${q_{t+1}}$ $={\Big{[}{{q_{t}}+{\alpha_{t}}\frac{{\partial{\mathcal{L}_{t}}({{x_{t}},q})}}{{\partial q}}\Big{|}{{}_{q={q_{t}}}}\Big{.}}\Big{]}_{+}},$ (12a)
${x_{t+1}}$ $={\mathcal{P}_{\mathbb{X}}}\Big{(}{{x_{t}}-{\alpha_{t}}\frac{{\partial{\mathcal{L}_{t}}({x,{q_{t+1}}})}}{{\partial x}}\Big{|}{{}_{x={x_{t}}}}}\Big{)},$ (12b)
where ${{\alpha_{t}}}$ is the stepsize. To calculate cumulative constraint violation, the clipped constraint function ${[{{g_{t}}({{x_{t}}})}]_{+}}$ is directly used to replace the original constraint function ${{g_{t}}({{x_{t}}})}$ in (11) in [28]. However, that makes Slater’s condition ineffective. To deal with this dilemma, we still use the original constraint function ${{g_{t}}({{x_{t}}})}$ in (11) but maximize ${{\mathcal{L}_{t}}({{x_{t}},q})}$ over all $q\in\mathbb{R}_{+}^{m}$ to replace (12a), i.e.,
${q_{t+1}}=\mathop{\arg\max}\limits_{q\in\mathbb{R}_{+}^{m}}{\mathcal{L}_{t}}({{x_{t}},q})={\gamma_{t}}{[{{g_{t}}({{x_{t}}})}]_{+}},$ (13)
where the second equation holds due to $-2{\gamma_{t}}{g_{t}}({{x_{t}}})/-2={\gamma_{t}}{g_{t}}({{x_{t}}})$. The updating rule (13) is also adopted in [26, 27, 29].
To implement the updating rules in a distributed manner, we use ${x_{i,t}}$ to denote the local copy of the primal variable ${x_{t}}$, and rewrite the dual variable in an agent-wise manner, i.e., ${q_{t}}={\rm{col}}({{q_{1,t}},\cdot\cdot\cdot,{q_{n,t}}})$ with each ${q_{1,t}}\in\mathbb{R}_{+}^{{m_{i}}}$. Then, the updating rule (13) can be executed in an agent-wise manner as (15a). Note that in the bandit setting, the subgradients are unavailable and only the values of ${f_{i,t}}$ and ${g_{i,t}}$ at some potential points are privately revealed. Thus, we use the values of the local loss function ${f_{i,t}}$ at ${x_{i,t}}$ and ${x_{i,t}}+{\delta_{t}}{u_{i,t}}$ to estimate the subgradient $\partial{f_{i,t}}({{x_{i,t}}})$, and use the values of the local constraint function ${g_{i,t}}$ at ${x_{i,t}}$ and ${x_{i,t}}+{\delta_{t}}{u_{i,t}}$ to estimate the subgradient $\partial{{g_{i,t}}({{x_{i,t}}})}$, i.e.,
^$\hat{\partial}{f_{i,t}}({{x_{i,t}}})$ $=\frac{p}{{{\delta_{t}}}}\big{(}{{f_{i,t}}({{x_{i,t}}+{\delta_{t}}{u_{i,t}}})-{f_{i,t}}({{x_{i,t}}})}\big{)}{u_{i,t}}\in{\mathbb{R}^{p}},$
^$\hat{\partial}{{g_{i,t}}({{x_{i,t}}})}$ $=\frac{p}{{{\delta_{t}}}}{\big{(}{{{{g_{i,t}}({{x_{i,t}}+{\delta_{t}}{u_{i,t}}})}}-{{{g_{i,t}}({{x_{i,t}}})}}}\big{)}^{T}}\otimes{u_{i,t}}\in{\mathbb{R}^{p\times{m_{i}}}},$
where ${\delta_{t}}\in({0,r(\mathbb{X}){\xi_{t}}}]$ is an exploration parameter, ${r(\mathbb{X})}$ is a positive constant, ${\xi_{t}}\in({0,1})$ is a shrinkage coefficient, and ${u_{i,t}}\in{\mathbb{S}^{p}}$ is a uniformly distributed random vector. The idea follows the two-point stochastic subgradient estimator proposed in [12, 36], and is also adopted in [28, 35]. ${{\hat{\omega}}_{i,t+1}}$ defined in (15b) can be understood as an estimator for a portion of $\frac{{\partial{\mathcal{L}_{t}}({x,{q_{t+1}}})}}{{\partial x}}\Big{|}{{}_{x={x_{t}}}}$ that is available to agent $i$. Then, each ${z_{i,t+1}}$ updated by (15c) can be understood as a local estimate of ${x_{t+1}}$ updated by (12b). To estimate ${x_{t+1}}$ more accurately, each agent $i$ computes ${x_{i,t+1}}$ by the consensus protocol (14), which tracks the average $\frac{1}{n}\sum\nolimits_{i=1}^{n}{{z_{i,t+1}}}$. As a result, the distributed bandit online primal–dual algorithm is proposed, which is presented in pseudo-code as Algorithm 1.
Algorithm 1 Distributed Bandit Online Primal–Dual Algorithm
0:  constant ${r(\mathbb{X})}$, non-increasing sequences $\{{\alpha_{t}}\}\subseteq({0,+\infty})$, $\{{{\xi_{t}}}\}\subseteq({0,1})$, $\{{{\delta_{t}}}\}\subseteq({0,r(\mathbb{X}){\xi_{t}}}]$, and non-decreasing sequence $\{{\gamma_{t}}\}\subseteq({0,+\infty})$.
0:  ${z_{i,1}}\in{(1-\xi_{1})\mathbb{X}}$.
  for $t=1,\cdot\cdot\cdot,T-1$ do
     for $i=1,\cdot\cdot\cdot,n$ in parallel do
        Broadcast $z_{i,t}$ to $\mathcal{N}_{i}^{\text{out}}({{\mathcal{G}_{t}}})$ and receive $z_{j,t}$ from $j\in\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})$.
        Select
${x_{i,t}}$ $=\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{z_{j,t}}}.$ (14)
        Select vector ${u_{i,t}}\in{\mathbb{S}^{p}}$ independently and uniformly at random.
        Observe ${{f_{i,t}}({{x_{i,t}}})}$, ${{f_{i,t}}({{x_{i,t}}+{\delta_{t}}{u_{i,t}}})}$, ${{{{g_{i,t}}({{x_{i,t}}})}}}$, ${{{{g_{i,t}}({{x_{i,t}}+{\delta_{t}}{u_{i,t}}})}}}$.
        Update
${q_{i,t+1}}$ $={\gamma_{t}}{[{{g_{i,t}}({{x_{i,t}}})}]_{+}},$ (15a)
${\hat{\omega}_{i,t+1}}$ =^$=\hat{\partial}{f_{i,t}}({x_{i,t}})+\hat{\partial}{{g_{i,t}}({x_{i,t}})}{q_{i,t+1}},$ (15b)
${z_{i,t+1}}$ $={\mathcal{P}_{({1-{\xi_{t+1}}})\mathbb{X}}}({{x_{i,t}}-{\alpha_{t}}{{\hat{\omega}}_{i,t+1}}}).$ (15c)
     end for
  end for
  $\{x_{i,t}\}$.
III-B Performance Analysis for Convex Case
In this subsection, we establish network regret and cumulative constraint violation bounds for Algorithm 1 in the following theorems when local loss functions are convex. We first establish these bounds without Slater’s condition.
Theorem 1.
Suppose Assumptions 1–4 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with
${\alpha_{t}}=\frac{1}{{{t^{c}}}},{\gamma_{t}}=\frac{{{\gamma_{0}}}}{{{\alpha_{t}}}},{\xi_{t}}={\alpha_{t}},{\delta_{t}}=r(\mathbb{X}){\alpha_{t}},$ (16)
where $c\in({0,1})$, ${\gamma_{0}}\in\big{(}{0,\sigma/({4(p^{2}+1)G_{2}^{2}})}\big{]}$ are constants. Then, for any $T\in{\mathbb{N}_{+}}$,
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $=\mathcal{O}({T^{\max\{{c,1-c}\}}}),$ (17)
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}({T^{1-c/2}}).$ (18)
The proof and the explicit expressions of the right-hand sides of (17)–(18) are given in Appendix B.
Remark 1.
In Theorem 1, we show that Algorithm 1 establishes sublinear network regret and cumulative constraint violation bounds. The bounds (17) and (18) are the same as the state-of-the-art bounds established by the distributed bandit online algorithm in [28]. Note that the potential drawback of Algorithm 1 is that it uses ${G_{2}}$ to design the algorithm parameter ${\gamma_{0}}$. However, we do not use the assumption that local constraint functions are uniformly bounded while [28] uses it. The bounds (17) and (18) are also the same as the bounds established by the distributed online algorithm with full-information feedback in [29]. If setting $c=1/2$, they then generalize the results established in [14, 15, 25], even though the bandit online algorithms in [14, 15] are centralized and the more tolerant constraint violation metric is used, and in [25] local loss functions are quadratic and local constraint functions are linear, static and known in advance. Moreover, the bounds (17) and (18) improve the results established by the distributed online algorithm with one-point bandit feedback in [27], even though in [27] local constraint functions are static and known in advance.
Remark 2.
Different from the algorithm in [29] that directly uses the subgradients of local loss and constraint functions, Algorithm 1 is based on the two-point stochastic gradient estimators, which are unbiased subgradients of the uniformly smoothed versions of local loss and constraint functions. In the one-dimensional case ($p=1$), the intuition is readily seen that the expectations of the estimators ^$\hat{\partial}{f_{i,t}}({{x_{i,t}}})$ and ^$\hat{\partial}{{g_{i,t}}({{x_{i,t}}})}$ equal $\frac{1}{{2{\delta_{t}}}}({f_{i,t}}({x_{i,t}}+{\delta_{t}})-{f_{i,t}}({x_{i,t}}-{\delta_{t}}))$ and $\frac{1}{{2{\delta_{t}}}}({g_{i,t}}({x_{i,t}}+{\delta_{t}})-{g_{i,t}}({x_{i,t}}-{\delta_{t}}))$. We know that they indeed approximate the derivatives of ${f_{i,t}}$ and ${g_{i,t}}$ at ${x_{i,t}}$ if ${\delta_{t}}$ is infinitesimal. However, ${\delta_{t}}$ cannot be small enough, and thus there exist gaps between ^$\hat{\partial}{f_{i,t}}({{x_{i,t}}})$, ^$\hat{\partial}{{g_{i,t}}({{x_{i,t}}})}$ and their true subgradients. These gaps prevent us from simply using the uniformly bounds of the subgradients in Assumption 3 as done in [29] to bound the estimators, and thus cause nontrivial challenges for performance analysis. To cope with this challenge, we need to analyze some properties of the estimators such as the uniformly bounds of the estimators and the gaps between the local loss and constraint functions and the correspondingly uniformly smoothed functions. Moreover, Algorithm 1 updates the dual variables by using the projected gradient descent instead of the composite objective mirror descent used in [29]. Therefore, the proof of our Theorem 1 has significant differences compared to that of Theorem 1 in [29].
With Slater’s condition, we show that Algorithm 1 establishes a reduced network cumulative constraint violation bound than the bound established in (18) in the following theorem.
Theorem 2.
Suppose Assumptions 1–5 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with (16). Then, for any $T\in{\mathbb{N}_{+}}$,
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $=\mathcal{O}({T^{\max\{{c,1-c}\}}}),$ (19)
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}({T^{1-c}}).$ (20)
The proof and the explicit expressions of the right-hand sides of (19)–(20) are given in Appendix C.
Remark 3.
As pointed out in [26], it is an open problem how to establish a reduced cumulative constraint violation bound for online convex optimization. Such a bound is first established by the distributed online algorithm with full-information feedback in [29]. In bandit convex optimization, such a bound is still missing as the method that directly using the clipped constraint functions to establish network cumulative constraint violation bound as used in [28] makes Slater’s condition ineffective. Theorem 2 establishes the reduced cumulative constraint violation bound (20), which is the same as the result established in [29], and thus fills the gap.
Remark 4.
Slater’s condition plays an important role for establishing the reduced network cumulative constraint violation bound (20). We will provide an elucidation of why network cumulative constraint violation bounds can be reduced under Slater’s condition and why directly using clipped constraint functions as done in [28] makes Slater’s condition ineffective. For the first question, it is worth noting that (47) in Lemma 5 is very important to establish network cumulative constraint violation bounds. Without Slater’s condition, from (58), we have $\frac{{q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}}{{{\gamma_{t}}}}\geq 0$, and then we have $\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{4{\gamma_{0}}}}}}\leq{{\tilde{h}}_{T}}({\hat{y}})$ since (44) and (73) (i.e., $\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\frac{{q_{i,t+1}^{T}{g_{i,t}}(y)}}{{{\gamma_{t}}}}}}\leq 0$) hold. Therefore, we have the result in (74). Based on the result, we can get (66) in Lemma 6, and thus establish the network cumulative constraint violation bound $\mathcal{O}\big{(}{{{({T\mathbf{E}[{{\tilde{h}}_{T}}(\hat{y})})]}^{1/2}}}\big{)}$ for the general cases. With Slater’s condition, we have (106) (i.e., $\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\frac{{q_{i,t+1}^{T}{g_{i,t}}(x_{s})}}{{{\gamma_{t}}}}}}\leq-{\varsigma_{s}}\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\|{{{[{{g_{i,t}}({x_{i,t}})}]}_{+}}}\|}}$). Note that the result in (106) is tighter than that in (73). By using ${\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{4{\gamma_{0}}}}}\geq 0$, we have a new lower bound ${\varsigma_{s}}\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\|{{{[{{g_{i,t}}({x_{i,t}})}]}_{+}}}\|}}$ for ${{\tilde{h}}_{T}}({\hat{y}})$ to replace the lower bound $\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{4{\gamma_{0}}}}}}$. As a result, by using (67) in Lemma 6, we establish the network cumulative constraint violation bound $\mathcal{O}\big{(}\mathbf{E}[{{{\tilde{h}}_{T}}({\hat{y}})}]\big{)}$. By appropriately designing the stepsize sequence $\{{{\alpha_{t}}}\}$, we can guarantee $\mathcal{O}\big{(}\mathbf{E}[{{{\tilde{h}}_{T}}({\hat{y}})}]\big{)}=\mathbf{o}(T)$. Therefore, the intuition is readily seen that $\mathcal{O}\big{(}\mathbf{E}[{{{\tilde{h}}_{T}}({\hat{y}})}]\big{)}$ is in general smaller with respect to $T$ than $\mathcal{O}\big{(}{{{({T\mathbf{E}[{{\tilde{h}}_{T}}(\hat{y})})]}^{1/2}}}\big{)}$. Based on the above elucidation, we know that network cumulative constraint violation bounds can be reduced under Slater’s condition. For the second question, it should be pointed out that (106) is a key result to reduce network cumulative constraint violation bounds as discussed in the elucidation of the first question. However, if we directly use clipped constraint functions, the term $\frac{1}{n}\sum\nolimits_{i=1}^{n}{q_{i,t+1}^{T}{g_{i,t}}(y)}$ in (33) would be replaced by $\frac{1}{n}\sum\nolimits_{i=1}^{n}{q_{i,t+1}^{T}{{[{{g_{i,t}}(y)}]}_{+}}}$, and then the result in (106) would become $\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\frac{{q_{i,t+1}^{T}{{[{{g_{i,t}}({x_{s}})}]}_{+}}}}{{{\gamma_{t}}}}}}\geq 0$. Note that the property of Slater’s condition in (10) does not work to establish the lower bound for ${{\tilde{h}}_{T}}({\hat{y}})$, i.e., Slater’s condition is ineffective. Based on the above elucidation, we know that directly using clipped constraint functions makes Slater’s condition ineffective.
III-C Performance Analysis for Strongly Convex Case
In the subsection, we establish network regret and cumulative constraint violation bounds for Algorithm 1 in the following theorems when local loss functions are strongly convex.
Assumption 6.
For any $i\in[n]$ and $t\in{\mathbb{N}_{+}}$, $\{{{f_{i,t}}(x)}\}$ are strongly convex with convexity parameter $\mu>0$ over $\mathbb{X}$, i.e., for all $x,y\in\mathbb{X}$,
${f_{i,t}}(x)\geq{f_{i,t}}(y)+\langle{x-y,\partial{f_{i,t}}(y)}\rangle+\frac{\mu}{2}\|{x-y}\|^{2}.$ (21)
When the convex parameter $\mu$ is unknown, we use the natural vanishing stepsize sequence as in (16) of Theorem 1.
Theorem 3.
Suppose Assumptions 1–4 and 6 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with (16). Then, for any $T\in{\mathbb{N}_{+}}$,
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $=\mathcal{O}({T^{1-c}}),$ (22)
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}({T^{1-c/2}}).$ (23)
Moreover, if Assumptions 5 also holds, then
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}({T^{1-c}}).$ (24)
The proof and the explicit expressions of the right-hand sides of (22)–(24) are given in Appendix D.
Remark 5.
Theorem 3 shows that a reduced network regret bound (22) is established compare to the bounds (17) and (19) established in Theorems 1 and 2, respectively. The bounds (22) and (23) are the same as those established by the distributed bandit online algorithm in [28]. It is worth noting that Algorithm 1 establishes a reduced network cumulative constraint violation bound (24) under Slater’s condition than the bound (23) without Slater’s condition, while the algorithm in [28] does not achieve such a result.
When the convex parameter $\mu$ is known, we appropriately design the stepsize sequence in the following theorem.
Theorem 4.
Suppose Assumptions 1–4 and 6 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with
${\alpha_{t}}=\frac{1}{{\mu t}},{\gamma_{t}}=\frac{{{\gamma_{0}}}}{{{\alpha_{t}}}},{\xi_{t}}={\alpha_{t}},{\delta_{t}}=r(\mathbb{X}){\alpha_{t}},$ (25)
where ${\gamma_{0}}\in\big{(}{0,\sigma/({4(p^{2}+1)G_{2}^{2}})}\big{]}$ is a constant. Then, for any $T\in{\mathbb{N}_{+}}$,
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $=\mathcal{O}(\log(T)),$ (26)
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}(\sqrt{\log(T)T}).$ (27)
Moreover, if Assumptions 5 also holds, then
$\mathbf{E}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}(\log(T)).$ (28)
The proof and the explicit expressions of the right-hand sides of (26)–(28) are given in Appendix E.
Remark 6.
Theorem 4 shows that the reduced network regret bound (26) and network cumulative constraint violation bounds (27) and (28) compare to the bounds (22)–(24), respectively. It should be pointed out that the bound (28) is achieved for the first time in the literature. In addition, the bounds (26) and (27) improve the results established by the distributed online algorithm with one-point bandit feedback in [27], even though in [27] local constraint functions are static and known in advance.
IV NUMERICAL EXAMPLE
To evaluate the performance of Algorithm 1, we consider a distributed online linear regression problem with time-varying linear inequality constraints over a network of $n$ agents. At iteration $t$, the local loss and constraint functions are ${f_{i,t}}(x)=\frac{1}{2}{({{A_{i,t}}x-{\vartheta_{i,t}}})^{2}}$ and ${g_{i,t}}(x)={B_{i,t}}x-{b_{i,t}}$, respectively, where each component of ${A_{i,t}}\in{\mathbb{R}^{{q_{i}}\times p}}$ is randomly generated from the uniform distribution in the interval $[{-1,1}]$, ${\vartheta_{i,t}}={A_{i,t}}{\mathbf{1}_{p}}+{\zeta_{i,t}}$ with ${\vartheta_{i,t}}\in{\mathbb{R}^{{q_{i}}}}$ and ${\zeta_{i,t}}$ being a standard normal random vector, each component of ${B_{i,t}}\in{\mathbb{R}^{{m_{i}}\times p}}$ is randomly generated from the uniform distribution in the interval $[{0,2}]$, each component of ${b_{i,t}}\in{\mathbb{R}^{{m_{i}}}}$ is randomly generated from the uniform distribution in the interval $[{b,b+1}]$ with $b>0$. Note that $b>0$ guarantees Slater’s condition holds. We use an time-varying undirected graph to model the communication topology. Specifically, at each iteration $t$, the graph is first randomly generated where the probability of any two agents being connected is $\rho$. Then, to make sure that Assumption 4 is satisfied, we add edges $({i,i+1})$ for $i\in[24]$ when $t\in\{{4c+1}\}$, edges $({i,i+1})$ for $i\in[25,49]$ when $t\in\{{4c+2}\}$, edges $({i,i+1})$ for $i\in[50,74]$ when $t\in\{{4c+3}\}$, edges $({i,i+1})$ for $i\in[75,99]$ when $t\in\{{4c+4}\}$ for $c=\{0,1,\cdot\cdot\cdot\}$. Moreover, let ${[{{W_{t}}}]_{ij}}=\frac{1}{n}$ if $({j,i})\in{\mathcal{E}_{t}}$ and ${[{{W_{t}}}]_{ii}}=1-\sum\nolimits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}$.
In this paper, without Slater’s condition, we show that Algorithm 1 establishes the same network regret and cumulative constraint violation bounds as those in [28]. More importantly, with Slater’s condition, we show that Algorithm 1 establishes the reduced network cumulative constraint violation bounds, which is significant results not found in existing literature. To verify our theoretical results, we compare Algorithm 1 with the distributed online algorithm with two-point bandit feedback in [28] that uses the cumulative constraint violation metric but does not consider Slater’s condition, and the distributed online algorithm with full-information feedback in [29] that uses the cumulative constraint violation metric and consider Slater’s condition. We set $n=100$, ${q_{i}}=4$, $p=10$, ${m_{i}}=2$, $\mathbb{X}={[{-5,5}]^{p}}$, $b=0.01$, and $\rho=0.1$. The inputs of the algorithms are listed in TABLE II.
TABLE II: Input of algorithms.
Algorithms Inputs
Algorithm 1 in this paper ${\alpha_{t}}=1/t,{\gamma_{t}}=0.15/{\alpha_{t}},{\xi_{t}}=1/t,{\delta_{t}}=0.01/t$
Algorithm 2 in [28] ${\alpha_{t}}=1/{t},{\beta_{t}}=1/{t^{0.5}},{\gamma_{t}}=1/{t^{0.5}},{\xi_{t}}=1/({t+1}),{\delta_{t}}=0.01/({t+1})$
Algorithm 1 in [29] ${\alpha_{t}}=2/t,{\gamma_{t}}=0.15/{\alpha_{t}},\psi(x)={\|x\|^{2}}$
Figure 1: Evolutions of $\frac{1}{n}\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}}$.
Figure 2: Evolutions of $\frac{1}{n}\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}$.
Figs. 1 and 2 illustrate the evolutions of the cumulative loss $\frac{1}{n}\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}}$ and the cumulative constraint violation $\frac{1}{n}\sum\nolimits_{i=1}^{n}{\sum\nolimits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}$, respectively. Fig. 1 demonstrates that our Algorithm 1 has almost the same accumulated loss as that of Algorithm 2 in [28], but has slightly larger accumulated loss than that of Algorithm 1 in [29]. That is reasonable since Algorithm 1 in [29] directly uses the subgradients of local loss and constraint functions while our Algorithm 1 uses two-point stochastic estimators to approximate these subgradients. Fig. 2 demonstrates that our Algorithm 1 has significantly smaller cumulative constraint violation than that of Algorithm 2 in [28], which are consistent with the theoretical results in Theorem 4. The key reason is that Slater’s condition remains effective in our Algorithm 1 but becomes ineffective in Algorithm 2 in [28].
V CONCLUSIONS
This paper studied the distributed bandit convex optimization problem with time-varying inequality constraints. We proposed a new distributed bandit online primal–dual algorithm, and established network regret and cumulative constraint violation bounds for convex and strongly convex cases, respectively. Without Slater’s condition, the bounds were the same as the state-of-the-art those in the literature. With Slater’s condition, the network cumulative constraint violation bounds were reduced. In the future, we will explore the scenario where communication resources are limited, and investigate distributed bandit online algorithms with compressed communication.
A. Useful Lemmas
Some preliminary results are given in this subsection.
We first provide some results on the regularized Bregman projection in the following lemma.
Lemma 1.
Let $\mathbb{K}$ be a nonempty closed convex subset of ${\mathbb{R}^{p}}$ and let $a$ and $b$ be two vectors in ${\mathbb{R}^{p}}$. If ${x_{c}}={\mathcal{P}_{\mathbb{K}}}({b-a})$, then
$2\langle{{x_{c}}-y,a}\rangle\leq{\|{y-b}\|^{2}}-{\|{y-{x_{c}}}\|^{2}}-{\|{{x_{c}}-b}\|^{2}},\forall y\in\mathbb{K}.$ (29)
In addition, If $\Phi(y)={\|{b-y}\|^{2}}+\langle{a,y}\rangle$, then we know $\Phi$ is a strongly convex function with convexity parameter $\sigma>0$ and ${x_{c}}=\mathop{\arg\min}\limits_{y\in\mathbb{K}}\Phi(y)$. We can obtain
$\|{{x_{c}}-b}\|\leq\frac{{\|a\|}}{\sigma}.$ (30)
Proof.
We know (29) holds from Lemma 3 in [35]. From the optimality condition, we have
${\|{b-{x_{c}}}\|^{2}}+\langle{a,{x_{c}}}\rangle\leq{\|{b-b}\|^{2}}+2\langle{a,b}\rangle-\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}.$
From ${\|{b-{x_{c}}}\|^{2}}\geq\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}$, and rearranging the above inequality give
$\sigma{\|{b-{x_{c}}}\|^{2}}\leq{\|{b-{x_{c}}}\|^{2}}+\frac{\sigma}{2}{\|{b-{x_{c}}}\|^{2}}\leq\langle{a,b-{x_{c}}}\rangle.$
Thus, we have
$\sigma{\|{b-{x_{c}}}\|^{2}}\leq\|a\|\|{b-{x_{c}}}\|.$
From the above inequality and $\sigma>0$, we can obtain (30). ∎
We then present some properties of the subgradient estimators $\partial{{\hat{f}}_{i,t}}$ and $\partial{{{{\hat{g}}_{i,t}}}}$ in the following lemma.
Lemma 2.
(Lemma 8 in [28]) If Assumption 3 holds. Then, ${{\hat{f}}_{i,t}}(x)$ and ${{{{\hat{g}}_{i,t}}(x)}}$ are convex on $({1-{\xi_{t}}})\mathbb{X}$. If ${{f}_{i,t}}(x)$ and ${{{{g}_{i,t}}(x)}}$ are strongly convex with convexity parameter $\mu>0$ over $\mathbb{X}$, Then, ${{\hat{f}}_{i,t}}(x)$ and ${{{{\hat{g}}_{i,t}}(x)}}$ are strongly convex with convexity parameter $\mu>0$ over $({1-{\xi_{t}}})\mathbb{X}$. Moreover, for any $i\in[n]$, $t\in{\mathbb{N}_{+}}$, $x\in({1-{\xi_{t}}})\mathbb{X}$, $q\in\mathbb{R}_{+}^{{m_{i}}}$,
$\partial{{\hat{f}}_{i,t}}(x)={\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\hat{\partial}{f_{i,t}}(x)}],$ (31a)
${f_{i,t}}(x)\leq{{\hat{f}}_{i,t}}(x)\leq{f_{i,t}}(x)+{G_{1}}{\delta_{t}},$ (31b)
^$\|{\hat{\partial}{f_{i,t}}(x)}\|\leq p{G_{1}},$ (31c)
$\partial{{{{\hat{g}}_{i,t}}(x)}}={\mathbf{E}_{{\mathfrak{U}_{t}}}}\big{[}{\hat{\partial}{{{g_{i,t}}(x)}}}\big{]},$ (31d)
${q^{T}}{{{g_{i,t}}(x)}}\leq{q^{T}}{{{{\hat{g}}_{i,t}}(x)}}\leq{q^{T}}{{{g_{i,t}}(x)}}+{G_{2}}{\delta_{t}}\|q\|,$ (31e)
^$\|{\hat{\partial}{{{g_{i,t}}(x)}}}\|\leq p{G_{2}},$ (31f)
where ${{\hat{f}}_{i,t}}(x)={\mathbf{E}_{v\in{\mathbb{B}^{p}}}}[{{f_{i,t}}({x+{\delta_{t}}v})}]$ and ${{{{\hat{g}}_{i,t}}(x)}}={\mathbf{E}_{v\in{\mathbb{B}^{p}}}}\big{[}{{{{g_{i,t}}({x+{\delta_{t}}v})}}}\big{]}$ with $v$ being chosen uniformly at random, and ${\mathfrak{U}_{t}}$ is the $\sigma$-algebra induced by the independent and identically distributed variables ${u_{1,t}},\cdot\cdot\cdot,{u_{n,t}}$.
We next quantify the disagreement among the local temporary primal variables $\{{{z_{i,t}}}\}$.
Lemma 3.
(Lemma 4 in [28]) If Assumption 4 holds. For all $i\in[n]$ and $t\in{\mathbb{N}_{+}}$, ${z_{i,t}}$ generated by Algorithm 1 satisfy
$\|{{z_{i,t}}-{{\bar{z}}_{t}}}\|$ $\leq\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{z_{j,1}}}\|}+\frac{1}{n}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t-1}^{z}}\|}+\|{\varepsilon_{i,t-1}^{z}}\|+\tau\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|},$ (32)
where ${{\bar{z}}_{t}}=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{z_{j,t}}}$, $\tau={({1-\omega/4{n^{2}}})^{-2}}>1$, $\lambda={({1-\omega/4{n^{2}}})^{1/B}}\in({0,1})$, and $\varepsilon_{i,t-1}^{z}={z_{i,{t}}}-{x_{i,t-1}}$.
We finally analyze regret at one iteration.
Lemma 4.
Suppose Assumptions 1–3 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 and $\{{{y_{t}}}\}$ be an arbitrary sequence in $\mathbb{X}$, then
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}+\frac{1}{n}\sum\limits_{i=1}^{n}{\big{(}{{f_{i,t}}({{x_{i,t}}})-{f_{i,t}}(y)}\big{)}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{q_{i,t+1}^{T}{g_{i,t}}(y)}+\frac{1}{n}\sum\limits_{i=1}^{n}{{\Delta_{i,t}}({\hat{y}})}+\frac{{{{\tilde{\Delta}}_{t}}}}{n}$
$+\frac{1}{n}\sum\limits_{i=1}^{n}{{G_{2}}}\big{(}{R(\mathbb{X}){\xi_{t}}+{\delta_{t}}}\big{)}\|{{q_{i,t+1}}}\|+{G_{1}}R(\mathbb{X}){\xi_{t}}+{G_{1}}{\delta_{t}},$ (33)
where
${\Delta_{i,t}}({\hat{y}})=\frac{1}{{{2\alpha_{t}}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}\big{[}{{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}-{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}}\big{]},$
${{\tilde{\Delta}}_{t}}=\sum\limits_{i=1}^{n}{({p{G_{1}}+p{G_{2}}\|{{q_{i,t+1}}}\|}){\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]}-\sum\limits_{i=1}^{n}{\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{2{\alpha_{t}}}}}.$
Proof.
From Assumption 3, we have
${f_{i,t}}(y)$ $\geq{f_{i,t}}(x)+\langle{\partial{f_{i,t}}(x),y-x}\rangle,\forall x,y\in\mathbb{X},$ (34)
${g_{i,t}}(y)$ $\geq{g_{i,t}}(x)+\partial{g_{i,t}}(x)({y-x}),\forall x,y\in\mathbb{X}.$ (35)
From (6), (9a), (31b) and $\hat{y}=({1-{\xi_{t}}})y$ for any $t\in{\mathbb{N}_{+}}$, we have
${{\hat{f}}_{i,t}}({\hat{y}})-{f_{i,t}}(y)$ $={f_{i,t}}({\hat{y}})-{f_{i,t}}(y)+{{\hat{f}}_{i,t}}({\hat{y}})-{f_{i,t}}({\hat{y}})$
$\leq{G_{1}}\|{\hat{y}-y}\|+{{\hat{f}}_{i,t}}({\hat{y}})-{f_{i,t}}({\hat{y}})$
$\leq{G_{1}}R(\mathbb{X}){\xi_{t}}+{G_{1}}{\delta_{t}}.$ (36)
We have
$q_{i,t+1}^{T}{{\hat{g}}_{i,t}}({\hat{y}})$ $\leq q_{i,t+1}^{T}{g_{i,t}}({\hat{y}})+{G_{2}}{\delta_{t}}\|{{q_{i,t+1}}}\|$
$=q_{i,t+1}^{T}\big{(}{{g_{i,t}}(y)+{g_{i,t}}({\hat{y}})-{g_{i,t}}(y)}\big{)}+{G_{2}}{\delta_{t}}\|{{q_{i,t+1}}}\|$
$\leq q_{i,t+1}^{T}{g_{i,t}}(y)+{G_{2}}\|{\hat{y}-y}\|\|{{q_{i,t+1}}}\|+{G_{2}}{\delta_{t}}\|{{q_{i,t+1}}}\|$
$\leq q_{i,t+1}^{T}{g_{i,t}}(y)+{G_{2}}\big{(}{R(\mathbb{X}){\xi_{t}}+{\delta_{t}}}\big{)}\|{{q_{i,t+1}}}\|,$ (37)
where the first inequality holds due to (31e); the second inequality holds due to (8b) and (35); and the last inequality holds due to (6) and $\hat{y}=\left({1-{\xi_{t}}}\right)y$.
We have
${{\hat{f}}_{i,t}}({{x_{i,t}}})-{{\hat{f}}_{i,t}}({\hat{y}})$ $\leq\langle{\partial{{\hat{f}}_{i,t}}({{x_{i,t}}}),{x_{i,t}}-\hat{y}}\rangle$
$=\langle{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\hat{\partial}{f_{i,t}}({{x_{i,t}}})}],{x_{i,t}}-\hat{y}}\rangle$
$={\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\langle{\hat{\partial}{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-\hat{y}}\rangle}]$
$={\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\langle{\hat{\partial}{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-z_{i,t+1}}\rangle+\langle{\hat{\partial}{f_{i,t}}({{x_{i,t}}}),{z_{i,t+1}}-\hat{y}}\rangle}]$
$\leq{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{p{G_{1}}\|{\varepsilon_{i,t}^{z}}\|+\langle{\hat{\partial}{f_{i,t}}({{x_{i,t}}}),{z_{i,t+1}}-\hat{y}}\rangle}],$ (38)
where the first inequality holds since ${x_{i,t}}$, $\hat{y}\in({1-{\xi_{t}}})\mathbb{X}$ and ${f_{i,t}}(x)$ is convex on $({1-{\xi_{t}}})\mathbb{X}$; the first equality holds due to (31a); the second equality holds since ${x_{i,t}}$ and $\hat{y}$ are independent of ${\mathfrak{U}_{t}}$; and the last inequality holds due to (31c).
From (15b), we have
^$\langle{\hat{\partial}{f_{i,t}}({{x_{i,t}}}),{z_{i,t+1}}-\hat{y}}\rangle$ $=\langle{{{\hat{\omega}}_{i,t+1}},{z_{i,t+1}}-\hat{y}}\rangle+\langle{\hat{\partial}{g_{i,t}}({{x_{i,t}}}){q_{i,t+1}},\hat{y}-{z_{i,t+1}}}\rangle$
$=\langle{{{\hat{\omega}}_{i,t+1}},{z_{i,t+1}}-\hat{y}}\rangle+\langle{\hat{\partial}{g_{i,t}}({{x_{i,t}}}){q_{i,t+1}},\hat{y}-{x_{i,t}}}\rangle$
+^$\;\;\;\;+\langle{\hat{\partial}{g_{i,t}}({{x_{i,t}}}){q_{i,t+1}},{x_{i,t}}-{z_{i,t+1}}}\rangle.$ (39)
We have
${\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\langle{\hat{\partial}{g_{i,t}}({{x_{i,t}}}){q_{i,t+1}},\hat{y}-{x_{i,t}}}\rangle}]$ $=\langle{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\hat{\partial}{g_{i,t}}({{x_{i,t}}})}]{q_{i,t+1}},\hat{y}-{x_{i,t}}}\rangle$
$=\langle{\partial{{\hat{g}}_{i,t}}({{x_{i,t}}}){q_{i,t+1}},\hat{y}-{x_{i,t}}}\rangle$
$\leq q_{i,t+1}^{T}{{\hat{g}}_{i,t}}({\hat{y}})-q_{i,t+1}^{T}{{\hat{g}}_{i,t}}({{x_{i,t}}})$
$\leq q_{i,t+1}^{T}{g_{i,t}}(y)-q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})+{G_{2}}\big{(}{R(\mathbb{X}){\xi_{t}}+{\delta_{t}}}\big{)}\|{{q_{i,t+1}}}\|,$ (40)
where the first equality holds since ${x_{i,t}}$, $q_{i,t+1}$ and $\hat{y}$ are independent of ${\mathfrak{U}_{t}}$; the last equality holds due to (31d); the first inequality holds since ${q_{i,t}}\geq{\mathbf{0}_{{m_{i}}}}$, ${x_{i,t}},\hat{y}\in({1-{\xi_{t}}})\mathbb{X}$ and ${{\hat{g}}_{i,t}}(x)$ is convex on $({1-{\xi_{t}}})\mathbb{X}$ as shown in Lemma 2; and the last inequality holds due to (31e) and (37).
From the Cauchy-Schwarz inequality and (8b), we have
^$\big{\langle}{{{\hat{\partial}{g_{i,t}}({{x_{i,t}}})}}{q_{i,t+1}},{x_{i,t}}-{z_{i,t+1}}}\big{\rangle}\leq p{G_{2}}\|{{q_{i,t+1}}}\|\|{\varepsilon_{i,t}^{z}}\|.$ (41)
Applying (29) to update (15c), we have
$\;\;\;\;\;\langle{{\hat{\omega}_{i,t+1}},{z_{i,t+1}}-\hat{y}}\rangle$
$\leq\frac{1}{{2{\alpha_{t}}}}\big{(}{{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}-{{\|{\hat{y}-{z_{i,t+1}}}\|}^{2}}-\|{\varepsilon_{i,t}^{z}}\|^{2}}\big{)}$
$=\frac{1}{{2{\alpha_{t}}}}\big{(}{{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}-{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}+{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}-{{\|{\hat{y}-{z_{i,t+1}}}\|}^{2}}-\|{\varepsilon_{i,t}^{z}}\|^{2}}\big{)}$
$={\Delta_{i,t}}({\hat{y}})+\frac{1}{{2{\alpha_{t}}}}\Big{(}{{\Big{\|}{\hat{y}-\sum\limits_{j=1}^{n}{{{[{{W_{t+1}}}]}_{ij}}{z_{j,t+1}}}}\Big{\|}^{2}}-{{\|{\hat{y}-{z_{i,t+1}}}\|}^{2}}-\|{\varepsilon_{i,t}^{z}}\|^{2}}\Big{)}$
$\leq{\Delta_{i,t}}({\hat{y}})+\frac{1}{{2{\alpha_{t}}}}\Big{(}{\sum\limits_{j=1}^{n}{{{[{{W_{t+1}}}]}_{ij}}{{\|{\hat{y}-{z_{j,t+1}}}\|}^{2}}}-{{\|{\hat{y}-{z_{i,t+1}}}\|}^{2}}-\|{\varepsilon_{i,t}^{z}}\|^{2}}\Big{)},$ (42)
where the second equality holds due to (14); and the last inequality holds since ${W_{t+1}}$ is doubly stochastic and ${\|\cdot\|^{2}}$ are convex.
Summing (31b), (36), (38)–(42) over $i\in[n]$, dividing by $n$, using $\sum\nolimits_{i=1}^{n}{{{[{{W_{t}}}]}_{ij}}=1},\forall t\in{\mathbb{N}_{+}}$, and rearranging terms yields (33). ∎
Finally, we bound local regret and (squared) cumulative constraint violation, the accumulated (squared) consensus error, and the changes caused by composite mirror descent in the following.
Lemma 5.
Suppose Assumptions 1–2 and 4–6 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with ${\gamma_{t}}={\gamma_{0}}/{\alpha_{t}}$, where ${\gamma_{0}}\in\big{(}{0,1/({4(p^{2}+1)G_{2}^{2}})}\big{]}$ is a constant. Then, for any $T\in{\mathbb{N}_{+}}$,
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\big{(}{{f_{i,t}}({{x_{i,t}}})-{f_{i,t}}(y)+\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{4{\alpha_{t}}}}}\big{)}}}\leq{m_{T}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\Delta_{i,t}}({\hat{y}})}},$ (43)
$\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{1}{2}}}\Big{(}{\frac{{q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}}{{{\gamma_{t}}}}+\frac{{\ {\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{2{\gamma_{0}}}}}\Big{)}\leq{h_{T}}(y)+{{\tilde{h}}_{T}}({\hat{y}}),$ (44)
$\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{x_{i,t}}-{x_{j,t}}}\|}}}\leq n{\tilde{\varepsilon}_{1}}+{\tilde{\varepsilon}_{2}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\|{\varepsilon_{i,t}^{z}}\|}},$ (45)
$\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{{\|{{x_{i,t}}-{x_{j,t}}}\|}^{2}}}}}\leq{{\tilde{\varepsilon}}_{3}}+{{\tilde{\varepsilon}}_{4}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}}},$ (46)
${\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]\leq\frac{1}{\sigma}\big{(}{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}.$ (47)
where
${m_{T}}=\sum\limits_{t=1}^{T}{{2{p^{2}{G_{1}^{2}}}{\alpha_{t}}}}+\sum\limits_{t=1}^{T}{\frac{{R{{(\mathbb{X})}^{2}}\xi_{t}^{2}}}{{4{\alpha_{t}}}}}+\sum\limits_{t=1}^{T}{\frac{{\delta_{t}^{2}}}{{4{\alpha_{t}}}}}+\sum\limits_{t=1}^{T}{{G_{1}}R(\mathbb{X}){\xi_{t}}}+\sum\limits_{t=1}^{T}{{G_{1}}{\delta_{t}}},$
${h_{T}}(y)=\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{q_{i,t+1}^{T}{g_{i,t}}(y)}}{{{\gamma_{t}}}}}},$
${{\tilde{h}}_{T}}({\hat{y}})=\sum\limits_{i=1}^{n}\frac{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}}]}{{{2\gamma_{0}}}}+\sum\limits_{t=1}^{T}{\frac{{2nF}}{{{\gamma_{t}}}}}+\sum\limits_{t=1}^{T}{\frac{{2n{p^{2}{G_{1}^{2}}}{\gamma_{0}}}}{{\gamma_{t}^{2}}}}+\sum\limits_{t=1}^{T}{\frac{{nR{{(\mathbb{X})}^{2}}\xi_{t}^{2}}}{{4{\gamma_{0}}}}}+\sum\limits_{t=1}^{T}{\frac{{n\delta_{t}^{2}}}{{4{\gamma_{0}}}}}$
$\;\;\;\;\;\;\;\;\;\;+\sum\limits_{t=1}^{T}{\frac{{n{G_{1}}R(\mathbb{X}){\xi_{t}}+n{G_{1}}{\delta_{t}}}}{{{\gamma_{t}}}}},{\tilde{\varepsilon}_{1}}=\frac{{2\tau}}{{\lambda({1-\lambda})}}\sum\limits_{i=1}^{n}{\|{{z_{i,1}}}\|},$
${{\tilde{\varepsilon}}_{2}}=\frac{{2({n\tau+2-2\lambda})}}{{1-\lambda}},{{\tilde{\varepsilon}}_{3}}=\frac{{16n{\tau^{2}}}}{{{\lambda^{2}}({1-{\lambda^{2}}})}}{\Big{(}{\sum\limits_{i=1}^{n}{\|{{z_{i,1}}}\|}}\Big{)}^{2}},{{\tilde{\varepsilon}}_{4}}=\frac{{16{n^{2}}{\tau^{2}}}}{{{{({1-\lambda})}^{2}}}}+32,$
and $\sigma>0$ is the uniformly bounded of convexity parameter of the function $\Phi$ as shown in Lemma 1 when $b={x_{i,t}}$ and $a={\alpha_{t}}{{\hat{\omega}}_{i,t+1}}$ for any $i\in[n]$ and $t\in{\mathbb{N}_{+}}$.
Proof.
We will show that (43)–(47) hold in the following, respectively.
(i) Noting that ${g_{i,t}}(y)\leq{\mathbf{0}_{{m_{i}}}}$, $\forall i\in[n]$, $\forall t\in{\mathbb{N}_{+}}$ when $y\in{\mathcal{X}_{T}}$, summing (33) over $t\in[T]$ gives
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\big{(}{{f_{i,t}}({{x_{i,t}}})-{f_{i,t}}({{y}})}\big{)}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{t=1}^{T}\Big{(}-q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})+\frac{1}{n}{{\tilde{\Delta}}_{t}}+{\Delta_{i,t}}({\hat{y}})$
$\;\;\;+{G_{2}}\big{(}{R(\mathbb{X}){\xi_{t}}+{\delta_{t}}}\big{)}\|{{q_{i,t+1}}}\|+{G_{1}}R(\mathbb{X}){\xi_{t}}+{G_{1}}{\delta_{t}}\Big{)}.$ (48)
We have
$\;\;\;\;\;\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{({p{G_{1}}+p{G_{2}}\|{{q_{i,t+1}}}\|}){\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]}}$
$\leq\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{({2{{p^{2}{G_{1}^{2}}}}{\alpha_{t}}}+{2p^{2}G_{2}^{2}{\alpha_{t}}{{\|{{q_{i,t+1}}}\|}^{2}}}+\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{4{\alpha_{t}}}})}}.$ (49)
We have
$\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{G_{2}}\big{(}{R(\mathbb{X}){\xi_{t}}+{\delta_{t}}}\big{)}\|{{q_{i,t+1}}}\|}}\leq\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\big{(}{{2G_{2}^{2}{\alpha_{t}}{{\|{{q_{i,t+1}}}\|}^{2}}}+\frac{{R{{(\mathbb{X})}^{2}}\xi_{t}^{2}}}{{4{\alpha_{t}}}}+\frac{{\delta_{t}^{2}}}{{4{\alpha_{t}}}}}\big{)}}}.$ (50)
From (15a), for all $t\in{\mathbb{N}_{+}}$, we have
$\|{{q_{i,t+1}}}\|={\gamma_{t}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|.$ (51)
Then, we have
${2(p^{2}+1)G_{2}^{2}{\alpha_{t}}{{\|{{q_{i,t+1}}}\|}^{2}}}-q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})=({{2(p^{2}+1)G_{2}^{2}{\gamma_{0}}}-1}){\gamma_{t}}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|^{2}}\leq 0$ (52)
where the equality holds due to the fact that $[b]_{+}^{T}b={\|{{{[b]}_{+}}}\|^{2}}$ for any vector $b$ and the inequality holds due to ${\gamma_{0}}\leq 1/({2(p^{2}+1)G_{2}^{2}})$.
Combining (48)–(50) and (52) yields (43).
(ii) From (7), we have
${f_{i,t}}(y)-{f_{i,t}}({{x_{i,t}}})\leq 2F,\forall y\in\mathbb{X}.$ (53)
Dividing (33) by ${{\gamma_{t}}}$, using (53), and summing over $t\in[T]$ gives
$\;\;\;\;\;\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}}{{{\gamma_{t}}}}}}$
$\;\;\leq{h_{T}}(y)+\sum\limits_{t=1}^{T}{\frac{{2nF}}{{{\gamma_{t}}}}}+\sum\limits_{t=1}^{T}{\frac{{{{\tilde{\Delta}}_{t}}}}{{{\gamma_{t}}}}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{\Delta_{i,t}}({\hat{y}})}}{{{\gamma_{t}}}}}}$
$\;\;+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{G_{2}}\big{(}{R(\mathbb{X}){\xi_{t}}+{\delta_{t}}}\big{)}\|{{q_{i,t+1}}}\|}}{{{\gamma_{t}}}}}}+\sum\limits_{t=1}^{T}{\frac{{n{G_{1}}R(\mathbb{X}){\xi_{t}}+n{G_{1}}{\delta_{t}}}}{{{\gamma_{t}}}}}.$ (54)
We have
$\;\;\;\;\;\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{({p{G_{1}}+p{G_{2}}\|{{q_{i,t+1}}}\|}){\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]}}{{{\gamma_{t}}}}}}$
$\;\;\leq\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\big{(}{\frac{{2{p^{2}{G_{1}^{2}}}{\gamma_{0}}}}{{\gamma_{t}^{2}}}+\frac{{2p^{2}G_{2}^{2}{\gamma_{0}}{{\|{{q_{i,t+1}}}\|}^{2}}}}{{\gamma_{t}^{2}}}+\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{4{\gamma_{0}}}}}\big{)}}}.$ (55)
We have
$\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{G_{2}}\big{(}{R(\mathbb{X}){\xi_{t}}+{\delta_{t}}}\big{)}\|{{q_{i,t+1}}}\|}}{{{\gamma_{t}}}}}}\leq\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\big{(}{\frac{{2G_{2}^{2}{\gamma_{0}}{{\|{{q_{i,t+1}}}\|}^{2}}}}{{\gamma_{t}^{2}}}+\frac{{R{{(\mathbb{X})}^{2}}\xi_{t}^{2}}}{{4{\gamma_{0}}}}+\frac{{\delta_{t}^{2}}}{{4{\gamma_{0}}}}}\big{)}}}.$ (56)
It follows from ${\gamma_{t}}{\alpha_{t}}={\gamma_{0}}$ and (6) that
$\sum\limits_{t=1}^{T}{\frac{{{\Delta_{i,t}}({\hat{y}})}}{{{\gamma_{t}}}}}=\sum\limits_{t=1}^{T}{\frac{1}{{{2\gamma_{0}}}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}-{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}}]\leq\frac{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}}]}{{{2\gamma_{0}}}}.$ (57)
From (15a) and the fact that $[b]_{+}^{T}b={\|{{{[b]}_{+}}}\|^{2}}$ for any vector $b$, we have
$\frac{{q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}}{{{\gamma_{t}}}}=[{{g_{i,t}}({{x_{i,t}}})}]_{+}^{T}{g_{i,t}}({{x_{i,t}}})={\|[{{g_{i,t}}({{x_{i,t}}})}]_{+}\|^{2}}.$ (58)
Combining (54)–(58) and (51), and using $1/2\geq 2(p^{2}+1)G_{2}^{2}{\gamma_{0}}$ yields (44).
[…]
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{x_{i,t}}-{x_{j,t}}}\|}}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\Big{\|}{\sum\limits_{l=1}^{n}{{{[{{W_{t}}}]}_{il}}{z_{l,t}}}-{{\bar{z}}_{t}}+{{\bar{z}}_{t}}-\sum\limits_{l=1}^{n}{{{[{{W_{t}}}]}_{j,l}}}{z_{l,t}}}\Big{\|}}}$
$\leq 2\sum\limits_{i=1}^{n}{\Big{\|}{\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}{z_{j,t}}}-{{\bar{z}}_{t}}}\Big{\|}}=2\sum\limits_{i=1}^{n}{\Big{\|}{\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}({{z_{j,t}}-{{\bar{z}}_{t}}})}}\Big{\|}}$
$\leq 2\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}\|{{z_{j,t}}-{{\bar{z}}_{t}}}\|}}=2\sum\limits_{i=1}^{n}{\|{{z_{i,t}}-{{\bar{z}}_{t}}}\|}.$ (59)
We have
$\sum\limits_{t=3}^{T}{\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}=\sum\limits_{t=1}^{T-2}{\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t}^{z}}\|}}\sum\limits_{s=0}^{T-t-2}{{\lambda^{s}}}\leq\frac{1}{{1-\lambda}}\sum\limits_{t=1}^{T-2}{\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t}^{z}}\|}}.$ (60)
We have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{x_{i,t}}-{x_{j,t}}}\|}}}$
$\leq\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{2\|{{z_{i,t}}-{{\bar{z}}_{t}}}\|}}$
$\leq\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{2\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{z_{j,1}}}\|}}}+\sum\limits_{t=2}^{T}{\sum\limits_{i=1}^{n}{2\Big{(}{\frac{1}{n}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t-1}^{z}}\|}+\|{\varepsilon_{i,t-1}^{z}}\|}\Big{)}}}$
$\;\;+\sum\limits_{t=3}^{T}{\sum\limits_{i=1}^{n}{2\tau}\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}}$
$\leq n{\tilde{\varepsilon}_{1}}+\sum\limits_{t=2}^{T}{\sum\limits_{i=1}^{n}{4\|{\varepsilon_{i,t-1}^{z}}\|}}+\frac{{2n\tau}}{{1-\lambda}}\sum\limits_{t=1}^{T-2}{\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t}^{z}}\|}},$ (61)
where the first inequality holds due to (59); the second inequality holds due to (32); and the last inequality holds due to (60). It follows from (61) that (45) holds.
(iv) Similar to the way to get (59), from (14), ${\|\cdot\|^{2}}$ is convex, and $\sum\nolimits_{i=1}^{n}{{{\left[{{W_{t}}}\right]}_{ij}}}=\sum\nolimits_{j=1}^{n}{{{\left[{{W_{t}}}\right]}_{ij}}}=1$, we have
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{{\|{{x_{i,t}}-{x_{j,t}}}\|}^{2}}}}\leq\sum\limits_{i=1}^{n}{4{{\|{{z_{i,t}}-{{\bar{z}}_{t}}}\|}^{2}}}.$ (62)
From (32), we have
     $\;\;\;\;\;4\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{{\|{{z_{i,t}}-{{\bar{z}}_{t}}}\|}^{2}}}}$
$\leq 4\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{z_{j,1}}}\|}+\|{\varepsilon_{i,t-1}^{z}}\|+\frac{1}{n}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t-1}^{z}}\|}+\tau\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}}}\Big{)}^{2}}}}$
$\leq 16\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\Big{(}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{z_{j,1}}}\|}}\Big{)}^{2}}+{{\|{\varepsilon_{i,t-1}^{z}}\|}^{2}}+{\Big{(}{\frac{1}{n}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t-1}^{z}}\|}}\Big{)}^{2}}+{\Big{(}{\tau\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}}}\Big{)}^{2}}}\Big{)}}}$
$\leq 16\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\Big{(}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{z_{j,1}}}\|}}\Big{)}^{2}}+{\|{\varepsilon_{i,t-1}^{z}}\|^{2}}+\frac{1}{n}\sum\limits_{j=1}^{n}{{\|{\varepsilon_{j,t-1}^{z}}\|^{2}}}+{\tau^{2}}\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}{\Big{(}{\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}}\Big{)}^{2}}}}\Big{)}}}$
$\leq 16\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\Big{(}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{z_{j,1}}}\|}}\Big{)}^{2}}+2{\|{\varepsilon_{i,t-1}^{z}}\|^{2}}+\frac{{n{\tau^{2}}}}{{1-\lambda}}\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}\sum\limits_{j=1}^{n}{{\|{\varepsilon_{j,s}^{z}}\|^{2}}}}\Big{)}}}$
$=16\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\Big{(}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{z_{j,1}}}\|}}\Big{)}^{2}}+2{\|{\varepsilon_{i,t-1}^{z}}\|^{2}}}\Big{)}}}+\frac{{16{n^{2}}{\tau^{2}}}}{{1-\lambda}}\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T-2}{{{\|{\varepsilon_{j,t}^{z}}\|}^{2}}\sum\limits_{s=0}^{T-t-2}{{\lambda^{s}}}}}$
$\leq{{\tilde{\varepsilon}}_{3}}+{{\tilde{\varepsilon}}_{4}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}}},$ (63)
where the third inequality holds due to the Hölder’s inequality. It follows form (62)–(63) that (46) holds.
(v) Applying (30) to update (15c) gives
${\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\varepsilon_{i,t}^{z}}\|]$ $={\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{{z_{i,t+1}}-{x_{i,t}}}\|]\leq\frac{{{\alpha_{t}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{{{\hat{\omega}}_{i,t+1}}}\|]}}{\sigma}$
$=\frac{{{\alpha_{t}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[\|{\hat{\partial}{f_{i,t}}({{x_{i,t}}})+\hat{\partial}{g_{i,t}}({{x_{i,t}}}){q_{i,t+1}}}\|]}}{\sigma}$
$\leq\frac{{{\alpha_{t}}}}{\sigma}\big{(}{{G_{1}}+{G_{2}}{\gamma_{t}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}$
$=\frac{1}{\sigma}\big{(}{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}.$ (64)
where the second equality holds due to (15b); the last inequality holds due to (8a), (8b) and (51); the last equality holds due to ${\gamma_{t}}{\alpha_{t}}={\gamma_{0}}$. It follows form (64) that (47) holds. ∎
Next, network regret and cumulative constraint violation bounds for the general cases are provided in the following lemma.
Lemma 6.
Under the same condition as stated in Lemma 5, for any $T\in{\mathbb{N}_{+}}$, it holds that
${\mathbf{E}}\Big{[}\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}}-\sum\limits_{t=1}^{T}{{f_{t}}(y)}\Big{]}\leq{\varepsilon_{1}}{{G_{1}}}+{n_{T}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\mathbf{E}}[{{\Delta_{i,t}}({\hat{y}})}]},$ (65)
${\mathbf{E}}\Big{[}\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}\Big{]}\leq\sqrt{2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}T+{\varepsilon_{2}}T{\mathbf{E}}[{{\tilde{h}}_{T}}({\hat{y}})}],$ (66)
${\mathbf{E}}\Big{[}\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}\Big{]}\leq n{G_{2}}{\tilde{\varepsilon}_{1}}+{\varepsilon_{3}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+{\varepsilon_{4}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}},$ (67)
where
${n_{T}}=\sum\limits_{t=1}^{T}{{\tilde{\varepsilon}_{2}^{2}{{{{G_{1}^{2}}}}}{\alpha_{t}}}}+{m_{T}},{\varepsilon_{2}}=\frac{{4\max\{{1,G_{2}^{2}{{\tilde{\varepsilon}}_{4}}}\}}}{{\min\{{1,\frac{\sigma}{{2{\gamma_{0}}}}}\}}},$
${\varepsilon_{3}}=\frac{{n{{G_{1}}}{G_{2}}{{\tilde{\varepsilon}}_{2}}}}{\sigma},{\varepsilon_{4}}=\frac{{G_{2}^{2}{{\tilde{\varepsilon}}_{2}}{\gamma_{0}}+\sigma}}{\sigma}.$
Proof.
We will show that (65)–(67) hold in the following, respectively.
(i) From ${f_{t}}(x)=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{f_{j,t}}(x)}$, we have
$\sum\limits_{i=1}^{n}{{f_{t}}({{x_{i,t}}})}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{f_{j,t}}({{x_{i,t}}})}}=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{f_{j,t}}({{x_{j,t}}})}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\big{(}{{f_{j,t}}({{x_{i,t}}})-{f_{j,t}}({{x_{j,t}}})}\big{)}}}$
$=\sum\limits_{i=1}^{n}{{f_{i,t}}({{x_{i,t}}})}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\big{(}{{f_{j,t}}({{x_{i,t}}})-{f_{j,t}}({{x_{j,t}}})}\big{)}}}$
$\leq\sum\limits_{i=1}^{n}{{f_{i,t}}({{x_{i,t}}})}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{G_{1}\|{{x_{i,t}}-{x_{j,t}}}\|}},$ (68)
where the inequality holds due to (9a).
From (45), we have
$\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{G_{1}}\|{{x_{i,t}}-{x_{j,t}}}\|}}}\leq n{\varepsilon_{1}}{{G_{1}}}+\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\Big{(}{{\tilde{\varepsilon}_{2}^{2}{{{{G_{1}^{2}}}}}{\alpha_{t}}}+\frac{{{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}}}{{4{\alpha_{t}}}}}\Big{)}}}.$ (69)
Combining (68)–(69) and (43) yields (65).
(ii) We have
${\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|^{2}}$ $={\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}-{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}+{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|^{2}}$
$\geq\frac{1}{2}{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|^{2}}-{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}-{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|^{2}}$
$\geq\frac{1}{2}{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|^{2}}-{\|{{g_{i,t}}({{x_{i,t}}})-{g_{i,t}}({{x_{j,t}}})}\|^{2}}$
$\geq\frac{1}{2}{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|^{2}}-G_{2}^{2}{\|{{x_{i,t}}-{x_{j,t}}}\|^{2}},$ (70)
where the second and the third inequalities hold due to the nonexpansive property of the projection ${[\cdot]_{+}}$ and (9b), respectively.
From ${g_{t}}(x)={\rm{col}}\big{(}{{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)}\big{)}$, we have
$\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}}}=\sum\limits_{t=1}^{T}{\sum\limits_{j=1}^{n}{{{\|{{{[{{g_{t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}}.$ (71)
Summing (70) over $i,j\in[n],t\in[T]$, dividing by $n$, and using (71) and (46) gives
$\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{j=1}^{n}{{{\|{{{[{{g_{t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}}\leq 2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}+\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{2\big{(}{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}+G_{2}^{2}{{\tilde{\varepsilon}}_{4}}{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}}\big{)}}}.$ (72)
From ${g_{i,t}}(y)\leq{\mathbf{0}_{{m_{i}}}},\forall i\in[n],\forall t\in{\mathbb{N}_{+}}$, when $y\in{\mathcal{X}_{T}}$, we have
${h_{T}}\left(y\right)\leq 0.$ (73)
Combining (72), (73) and (44) gives
${\mathbf{E}}\Big{[}\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}}\Big{]}\leq 2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}+{\varepsilon_{2}}{\mathbf{E}}[{{\tilde{h}}_{T}}({\hat{y}})],\forall y\in{\mathcal{X}_{T}}.$ (74)
We have
${\Big{(}{\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}}\Big{)}^{2}}\leq\frac{T}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}}.$ (75)
Combining (74) and (75) yields (66).
(iii) From (45) and (47), we have
${\mathbf{E}}\Big{[}\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{x_{i,t}}-{x_{j,t}}}\|}}}\Big{]}$ $\leq n{\tilde{\varepsilon}_{1}}+{{\tilde{\varepsilon}}_{2}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\|{\varepsilon_{i,t}^{z}}\|}]}$
$\leq n{\tilde{\varepsilon}_{1}}+\frac{{{{\tilde{\varepsilon}}_{2}}}}{\sigma}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\big{(}{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}\big{)}}}.$ (76)
From (9b), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{j,t}}})}]}_{+}}}\|}}\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}}}$
$=\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}+{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}-{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}}}$
$\leq\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\big{(}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|+{G_{2}}\|{{x_{i,t}}-{x_{j,t}}}\|}\big{)}}}}.$ (77)
Combining (76) and (77) yields (67). ∎
B. Proof of Theorem 1
We will show that (17) and (18) in Theorem 1 hold in the following, respectively.
(i) From (6), we have
$\;\;\sum\limits_{t=1}^{T}{\frac{1}{{{\alpha_{t}}}}}\big{(}{{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}-{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}}\big{)}$
$=\sum\limits_{t=1}^{T}{\Big{(}{\frac{1}{{{\alpha_{t-1}}}}{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}-\frac{1}{{{\alpha_{t}}}}{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}+\Big{(}{\frac{1}{{{\alpha_{t}}}}-\frac{1}{{{\alpha_{t-1}}}}}\Big{)}{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}}\Big{)}}$
$\leq\frac{1}{{{\alpha_{0}}}}{\|{\hat{y}-{x_{i,1}}}\|^{2}}-\frac{1}{{{\alpha_{T}}}}{\|{\hat{y}-{x_{i,T+1}}}\|^{2}}+\Big{(}{\frac{1}{{{\alpha_{T}}}}-\frac{1}{{{\alpha_{0}}}}}\Big{)}4R{(\mathbb{X})^{2}}\leq\frac{{4R{{(\mathbb{X})}^{2}}}}{{{\alpha_{T}}}}.$ (78)
From (16), we have
[…]
$\sum\limits_{t=1}^{T}{\frac{{\xi_{t}^{2}}}{{{\alpha_{t}}}}}$ $=\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{{{T^{1-c}}}}{{1-c}},$ (80)
$\sum\limits_{t=1}^{T}{\frac{{\delta_{t}^{2}}}{{{\alpha_{t}}}}}$ $=r{(\mathbb{X})^{2}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{{r{{(\mathbb{X})}^{2}}{T^{1-c}}}}{{1-c}},$ (81)
$\sum\limits_{t=1}^{T}{{\xi_{t}}}$ $=\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{{{T^{1-c}}}}{{1-c}},$ (82)
$\sum\limits_{t=1}^{T}{{\delta_{t}}}$ $=r{(\mathbb{X})}\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{{r{{(\mathbb{X})}}{T^{1-c}}}}{{1-c}},$ (83)
$\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{T}}}}$ $=2R{{(\mathbb{X})}^{2}}{T^{c}}.$ (84)
Denote
${x^{*}}=\mathop{\arg\min}\limits_{x\in{\mathcal{X}_{T}}}\sum\limits_{t=1}^{T}{{l_{t}}(x)}.$
Choosing $y={x^{*}}$, and combining (65) and (78)–(84) gives
${\mathbf{E}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\varepsilon_{1}}{{G_{1}}}+\frac{{{{{{G_{1}^{2}}}}}\tilde{\varepsilon}_{2}^{2}{T^{1-c}}}}{{({1-c})}}+\frac{{2{{({p^{2}{G_{1}^{2}}})}}{T^{1-c}}}}{{({1-c})}}+\frac{{R{{(\mathbb{X})}^{2}}{T^{1-c}}}}{{4({1-c})}}$
$\;\;+\frac{{r{(\mathbb{X})^{2}}{T^{1-c}}}}{{4({1-c})}}+\frac{{{{G_{1}}}R(\mathbb{X}){T^{1-c}}}}{{1-c}}+\frac{{{G_{1}}r(\mathbb{X}){T^{1-c}}}}{{1-c}}+2R{{(\mathbb{X})}^{2}}{T^{c}},$ (85)
which yields (17).
(ii) From (6), we have
$\sum\limits_{i=1}^{n}\frac{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}}]}{{{2\gamma_{0}}}}\leq\frac{{2nR{{(\mathbb{X})}^{2}}}}{{{\gamma_{0}}}}.$ (86)
From (16), we have
$\sum\limits_{t=1}^{T}{\frac{1}{{{\gamma_{t}}}}}=\sum\limits_{t=1}^{T}{\frac{{{\alpha_{t}}}}{{{\gamma_{0}}}}}\leq\frac{{{T^{1-c}}}}{{{\gamma_{0}}({1-c})}}.$ (87)
When $c\in({0,1/2})$, from (16), we have
$\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{{T^{1-2c}}}}{{1-2c}},$
$\sum\limits_{t=1}^{T}{\frac{1}{{\gamma_{t}^{2}}}}=\sum\limits_{t=1}^{T}{\frac{{\alpha_{t}^{2}}}{{\gamma_{0}^{2}}}}\leq\frac{{{T^{1-2c}}}}{{\gamma_{0}^{2}({1-2c})}},$ (88)
$\sum\limits_{t=1}^{T}{\xi_{t}^{2}}=\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{{T^{1-2c}}}}{{1-2c}},$ (89)
$\sum\limits_{t=1}^{T}{\delta_{t}^{2}}=r{(\mathbb{X})^{2}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{r{{(\mathbb{X})}^{2}}{T^{1-2c}}}}{{1-2c}},$ (90)
$\sum\limits_{t=1}^{T}{\frac{{{\xi_{t}}}}{{{\gamma_{t}}}}}=\frac{1}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{{T^{1-2c}}}}{{{\gamma_{0}}({1-2c})}},$ (91)
$\sum\limits_{t=1}^{T}{\frac{{{\delta_{t}}}}{{{\gamma_{t}}}}}=\frac{{r(\mathbb{X})}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{r(\mathbb{X}){T^{1-2c}}}}{{{\gamma_{0}}({1-2c})}}.$ (92)
Combining (66) and (86)–(92) yields
$\;\;\;\;\;{\mathbf{E}}\Big{[}{\Big{(}{\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}}\Big{)}^{2}}\Big{]}$
$\leq 2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}T+\frac{{2n{R{{(\mathbb{X})}^{2}}}{\varepsilon_{2}}T}}{{{\gamma_{0}}}}+\frac{{2nF{\varepsilon_{2}}{T^{2-c}}}}{{{\gamma_{0}}({1-c})}}+\frac{{2n{{p^{2}{G_{1}^{2}}}}{\varepsilon_{2}}{T^{2-2c}}}}{{{\gamma_{0}}({1-2c})}}+\frac{{nR{{(\mathbb{X})}^{2}}{\varepsilon_{2}}{T^{2-2c}}}}{{4{\gamma_{0}}({1-2c})}}$
$\;\;+\frac{{nr{{(\mathbb{X})}^{2}}{\varepsilon_{2}}{T^{2-2c}}}}{{4{\gamma_{0}}({1-2c})}}+\frac{{n{{G_{1}}}R(\mathbb{X}){\varepsilon_{2}}{T^{2-2c}}}}{{{\gamma_{0}}({1-2c})}}+\frac{{n{G_{1}}r(\mathbb{X}){\varepsilon_{2}}{T^{2-2c}}}}{{{\gamma_{0}}({1-2c})}}.$ (93)
When $c=1/2$, from (16), we have
$\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq 2\log(T),$
$\sum\limits_{t=1}^{T}{\frac{1}{{\gamma_{t}^{2}}}}=\sum\limits_{t=1}^{T}{\frac{{\alpha_{t}^{2}}}{{\gamma_{0}^{2}}}}\leq\frac{{2\log(T)}}{{\gamma_{0}^{2}}},$ (94)
$\sum\limits_{t=1}^{T}{\xi_{t}^{2}}=\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq 2\log(T),$ (95)
$\sum\limits_{t=1}^{T}{\delta_{t}^{2}}=r{(\mathbb{X})^{2}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq 2r{(\mathbb{X})^{2}}\log(T),$ (96)
$\sum\limits_{t=1}^{T}{\frac{{{\xi_{t}}}}{{{\gamma_{t}}}}}=\frac{1}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{2\log(T)}}{{{\gamma_{0}}}},$ (97)
$\sum\limits_{t=1}^{T}{\frac{{{\delta_{t}}}}{{{\gamma_{t}}}}}=\frac{{r(\mathbb{X})}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{2r(\mathbb{X})\log(T)}}{{{\gamma_{0}}}}.$ (98)
Combining (66), (86), (87), and (94)–(98) yields
$\;\;\;\;\;{\mathbf{E}}\Big{[}{\Big{(}{\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}}\Big{)}^{2}}\Big{]}$
$\leq 2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}T+\frac{{2nR{(\mathbb{X})^{2}}{\varepsilon_{2}}T}}{{{\gamma_{0}}}}+\frac{{2nF{\varepsilon_{2}}{T^{2-c}}}}{{{\gamma_{0}}({1-c})}}+\frac{{4n{{{p^{2}{G_{1}^{2}}}}}{\varepsilon_{2}}T\log(T)}}{{{\gamma_{0}}}}+\frac{{nR{{(\mathbb{X})}^{2}}{\varepsilon_{2}}T\log(T)}}{{2{\gamma_{0}}}}$
$\;\;+\frac{{nr{{(\mathbb{X})}^{2}}{\varepsilon_{2}}T\log(T)}}{{2{\gamma_{0}}}}+\frac{{2n{{G_{1}}}R(\mathbb{X}){\varepsilon_{2}}T\log(T)}}{{{\gamma_{0}}}}+\frac{{2n{G_{1}}r(\mathbb{X}){\varepsilon_{2}}T\log(T)}}{{{\gamma_{0}}}}.$ (99)
When $c=({1/2,1})$, from (16), there exists a constant $Q>0$ such that
$\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq Q,$
$\sum\limits_{t=1}^{T}{\frac{1}{{\gamma_{t}^{2}}}}=\sum\limits_{t=1}^{T}{\frac{{\alpha_{t}^{2}}}{{\gamma_{0}^{2}}}}\leq\frac{Q}{{\gamma_{0}^{2}}},$ (100)
$\sum\limits_{t=1}^{T}{\xi_{t}^{2}}=\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq Q,$ (101)
$\sum\limits_{t=1}^{T}{\delta_{t}^{2}}=r{(\mathbb{X})^{2}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq r{(\mathbb{X})^{2}}Q,$ (102)
$\sum\limits_{t=1}^{T}{\frac{{{\xi_{t}}}}{{{\gamma_{t}}}}}=\frac{1}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{Q}{{{\gamma_{0}}}},$ (103)
$\sum\limits_{t=1}^{T}{\frac{{{\delta_{t}}}}{{{\gamma_{t}}}}}=\frac{{r(\mathbb{X})}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{r(\mathbb{X})Q}}{{{\gamma_{0}}}}.$ (104)
Combining (66), (86), (87), and (100)–(104) yields
$\;\;\;\;\;{\mathbf{E}}\Big{[}{\Big{(}{\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}}\Big{)}^{2}}\Big{]}$
$\leq 2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}T+\frac{{2n{R{{(\mathbb{X})}^{2}}}{\varepsilon_{2}}T}}{{{\gamma_{0}}}}+\frac{{2nF{\varepsilon_{2}}{T^{2-c}}}}{{{\gamma_{0}}({1-c})}}+\frac{{2n{{{p^{2}{G_{1}^{2}}}}}{\varepsilon_{2}}QT}}{{{\gamma_{0}}}}+\frac{{nR{{(\mathbb{X})}^{2}}{\varepsilon_{2}}QT}}{{4{\gamma_{0}}}}$
$\;\;+\frac{{nr{{(\mathbb{X})}^{2}}{\varepsilon_{2}}QT}}{{4{\gamma_{0}}}}+\frac{{n{{G_{1}}}R(\mathbb{X}){\varepsilon_{2}}QT}}{{{\gamma_{0}}}}+\frac{{n{G_{1}}r(\mathbb{X}){\varepsilon_{2}}QT}}{{{\gamma_{0}}}}.$ (105)
It follows form (93), (99), and (105) that (18) holds.
C. Proof of Theorem 2
We will show that (19) and (20) in Theorem 2 hold in the following, respectively.
(i) From (85), we have (19).
(ii) When Slater’s condition holds, we have
${h_{T}}({{x_{s}}})$ $=\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{q_{i,t+1}^{T}{g_{i,t}}({{x_{s}}})}}{{{\gamma_{t}}}}}}=\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{[{{g_{i,t}}({{x_{i,t}}})}]_{+}^{T}{g_{i,t}}({{x_{s}}})}}$
$\leq-\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\varsigma_{s}}[{{g_{i,t}}({{x_{i,t}}})}]_{+}^{T}{1_{{m_{i}}}}}}=-{\varsigma_{s}}{\sum\limits_{i=1}^{n}\sum\limits_{t=1}^{T}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|_{1}}$
$\leq-{\varsigma_{s}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}},$ (106)
where the second equality holds due to (15a) and the first inequality holds due to (10).
Choosing $y={x_{s}}$ in (44), and using (15a) and (106) gives
${\varsigma_{s}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}}\leq{{\tilde{h}}_{T}}({{\hat{y}}}).$ (107)
When $c\in({0,1/2})$, combining (107), (16) and (86)–(92) gives
$\;\;\;\;\;{\mathbf{E}}\Big{[}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}}\Big{]}$
$\leq\frac{{2n{R{{(\mathbb{X})}^{2}}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2nF{T^{1-c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-c})}}+\frac{{2n{{{p^{2}{G_{1}^{2}}}}}{T^{1-2c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-2c})}}+\frac{{nR{{(\mathbb{X})}^{2}}{T^{1-2c}}}}{{4{\varsigma_{s}}{\gamma_{0}}({1-2c})}}+\frac{{nr{{(\mathbb{X})}^{2}}{T^{1-2c}}}}{{4{\varsigma_{s}}{\gamma_{0}}({1-2c})}}$
$\;\;+\frac{{n{{G_{1}}}R(\mathbb{X}){T^{1-2c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-2c})}}+\frac{{n{G_{1}}r(\mathbb{X}){T^{1-2c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-2c})}}.$ (108)
From (16), (67), (108), we have
$\;\;\;\;\;{\mathbf{E}}[{\rm{Net}\mbox{-}\rm{CCV}}(T)]$
$\leq n{G_{2}}{\tilde{\varepsilon}_{1}}+\frac{{{\varepsilon_{3}}{T^{1-c}}}}{{1-c}}+\frac{{2n{R{{(\mathbb{X})}^{2}}}{\varepsilon_{4}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2nF{\varepsilon_{4}}{T^{1-c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-c})}}+\frac{{2n{{{p^{2}{G_{1}^{2}}}}}{\varepsilon_{4}}{T^{1-2c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-2c})}}+\frac{{nR{{(\mathbb{X})}^{2}}{\varepsilon_{4}}{T^{1-2c}}}}{{4{\varsigma_{s}}{\gamma_{0}}({1-2c})}}$
$\;\;+\frac{{nr{{(\mathbb{X})}^{2}}{\varepsilon_{4}}{T^{1-2c}}}}{{4{\varsigma_{s}}{\gamma_{0}}({1-2c})}}+\frac{{n{{G_{1}}}R(\mathbb{X}){\varepsilon_{4}}{T^{1-2c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-2c})}}+\frac{{n{G_{1}}r(\mathbb{X}){\varepsilon_{4}}{T^{1-2c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-2c})}}.$ (109)
When $c=1/2$, combining (107), (16), (86), (87), and (94)–(98) gives
$\;\;\;\;\;{\mathbf{E}}\Big{[}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}}\Big{]}$
$\leq\frac{{2n{R{{(\mathbb{X})}^{2}}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2nF{T^{1-c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-c})}}+\frac{{4n{{{p^{2}{G_{1}^{2}}}}}\log(T)}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{nR{{(\mathbb{X})}^{2}}\log(T)}}{{2{\varsigma_{s}}{\gamma_{0}}}}+\frac{{nr{{(\mathbb{X})}^{2}}\log(T)}}{{2{\varsigma_{s}}{\gamma_{0}}}}$
$\;\;+\frac{{2n{{G_{1}}}R(\mathbb{X})\log(T)}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}r(\mathbb{X})\log(T)}}{{{\varsigma_{s}}{\gamma_{0}}}}.$ (110)
From (16), (67), (110), we have
$\;\;\;\;\;{\mathbf{E}}[{\rm{Net}\mbox{-}\rm{CCV}}(T)]$
$\leq n{G_{2}}{\tilde{\varepsilon}_{1}}+\frac{{{\varepsilon_{3}}{T^{1-c}}}}{{1-c}}+\frac{{2n{R{{(\mathbb{X})}^{2}}}{\varepsilon_{4}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2nF{\varepsilon_{4}}{T^{1-c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-c})}}+\frac{{4n{{{p^{2}{G_{1}^{2}}}}}{\varepsilon_{4}}\log(T)}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{nR{{(\mathbb{X})}^{2}}{\varepsilon_{4}}\log(T)}}{{2{\varsigma_{s}}{\gamma_{0}}}}$
$\;\;+\frac{{nr{{(\mathbb{X})}^{2}}{\varepsilon_{4}}\log(T)}}{{2{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{{G_{1}}}R(\mathbb{X}){\varepsilon_{4}}\log(T)}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}r(\mathbb{X}){\varepsilon_{4}}\log(T)}}{{{\varsigma_{s}}{\gamma_{0}}}}.$ (111)
When $c=({1/2,1})$, combining (107), (16), (86), (87), and (100)–(104) gives
$\;\;\;\;\;{\mathbf{E}}\Big{[}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}}\Big{]}$
$\leq\frac{{2n{R{{(\mathbb{X})}^{2}}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2nF{T^{1-c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-c})}}+\frac{{2n{{{p^{2}{G_{1}^{2}}}}}Q}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{nR{{(\mathbb{X})}^{2}}Q}}{{4{\varsigma_{s}}{\gamma_{0}}}}+\frac{{nr{{(\mathbb{X})}^{2}}Q}}{{4{\varsigma_{s}}{\gamma_{0}}}}+\frac{{n{{G_{1}}}R(\mathbb{X})Q}}{{{\varsigma_{s}}{\gamma_{0}}}}$
$\;\;+\frac{{n{G_{1}}r(\mathbb{X})Q}}{{{\varsigma_{s}}{\gamma_{0}}}}.$ (112)
From (16), (67), (112), we have
$\;\;\;\;\;{\mathbf{E}}[{\rm{Net}\mbox{-}\rm{CCV}}(T)]$
$\leq n{G_{2}}{\tilde{\varepsilon}_{1}}+\frac{{{\varepsilon_{3}}{T^{1-c}}}}{{1-c}}+\frac{{2n{R{{(\mathbb{X})}^{2}}}{\varepsilon_{4}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2nF{\varepsilon_{4}}{T^{1-c}}}}{{{\varsigma_{s}}{\gamma_{0}}({1-c})}}+\frac{{2n{{{p^{2}{G_{1}^{2}}}}}Q{\varepsilon_{4}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{nR{{(\mathbb{X})}^{2}}Q{\varepsilon_{4}}}}{{4{\varsigma_{s}}{\gamma_{0}}}}$
$\;\;+\frac{{nr{{(\mathbb{X})}^{2}}Q{\varepsilon_{4}}}}{{4{\varsigma_{s}}{\gamma_{0}}}}+\frac{{n{{G_{1}}}R(\mathbb{X})Q{\varepsilon_{4}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{n{G_{1}}r(\mathbb{X})Q{\varepsilon_{4}}}}{{{\varsigma_{s}}{\gamma_{0}}}}.$ (113)
It follows from (109), (111), and (113) that (20) holds.
D. Proof of Theorem 3
We will show that (22)–(24) in Theorem 3 hold in the following, respectively.
Since Assumption 6 holds, (38) can be replaced by
${\hat{f}_{i,t}}({{x_{i,t}}})-{\hat{f}_{i,t}}({\hat{y}})\leq{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{p{G_{1}}\|{\varepsilon_{i,t}^{z}}\|+\langle{\hat{\partial}{f_{i,t}}({{x_{i,t}}}),{z_{i,t+1}}-\hat{y}}\rangle-\frac{\mu}{2}{\|{\hat{y}-{x_{i,t}}}\|^{2}}}].$ (114)
Note that different from (38), there exists an extra term $\frac{\mu}{2}{\|{\hat{y}-{x_{i,t}}}\|^{2}}$ in (114), and thus (65) can be replaced by
$\;\;\;\;\;{\mathbf{E}}\Big{[}\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}}-\sum\limits_{t=1}^{T}{{f_{t}}(y)}\Big{]}$
$\leq{\varepsilon_{1}}{G_{1}}+{n_{T}}+{\mathbf{E}}[\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\big{(}{{\Delta_{i,t}}(\hat{y})-\frac{\mu}{2}{\|{\hat{y}-{x_{i,t}}}\|^{2}}}\big{)}}}].$ (115)
Moreover, (44) can be replaced by
$\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\frac{1}{2}}}\Big{(}{\frac{{q_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}}{{{\gamma_{t}}}}+\frac{{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}}{{2{\gamma_{0}}}}}\Big{)}\leq{h_{T}}(y)+{{\tilde{h}}_{T}}({\hat{y}})+{{\hat{h}}_{T}}({\hat{y}}),$ (116)
where
${{\hat{h}}_{T}}({\hat{y}})=-\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{\mu{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}}]}}{{{2\gamma_{t}}}}}}.$
As a result, (66) can be replaced by
${\mathbf{E}}\Big{[}\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}\Big{]}\leq\sqrt{2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}T+{\varepsilon_{2}}T{\mathbf{E}}\big{[}\big{(}{{{\tilde{h}}_{T}}({\hat{y}})+{{\hat{h}}_{T}}({\hat{y}})}\big{)}\big{]}}.$ (117)
(i) We have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\big{(}{{\Delta_{i,t}}({\hat{y}}){-\frac{\mu}{2}{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}}}\big{)}}}$
$=\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}}\Big{(}{{\frac{1}{{2{\alpha_{t}}}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}-{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}]-\frac{\mu}{2}{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}}\Big{)}$
$=\frac{1}{{2n}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\Big{(}{\frac{1}{{{\alpha_{t-1}}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}]-\frac{1}{{{\alpha_{t}}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\hat{y}-{x_{i,t+1}}}\|}^{2}}]+\Big{(}{\frac{1}{{{\alpha_{t}}}}-\frac{1}{{{\alpha_{t-1}}}}-\mu}\Big{)}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}]}\Big{)}}}$
$\leq\frac{1}{{2n}}\sum\limits_{i=1}^{n}{\Big{(}{\frac{1}{{{\alpha_{0}}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}]-\frac{1}{{{\alpha_{T}}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\hat{y}-{x_{i,T+1}}}\|}^{2}}]+\sum\limits_{t=1}^{T}{\Big{(}{\frac{1}{{{\alpha_{t}}}}-\frac{1}{{{\alpha_{t-1}}}}-\mu}\Big{)}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}}]}\Big{)}}.$ (118)
Denote
${\varepsilon_{5}}=\Big{\lceil}{{{\Big{(}{\frac{1}{\mu}}\Big{)}}^{\frac{1}{{1-c}}}}}\Big{\rceil}.$
From (16), we have
$\frac{1}{{{\alpha_{t+1}}}}-\frac{1}{{{\alpha_{t}}}}-\mu=\frac{{t+1}}{{{{({t+1})}^{1-c}}}}-\frac{t}{{{t^{1-c}}}}-\mu<\frac{1}{{{t^{1-c}}}}-\mu\leq 0,\forall t\geq{\varepsilon_{5}}.$ (119)
Choosing $y={x^{*}}\in{\mathcal{X}_{T}}$, and combining (6), (79)–(83), (115), and (118)–(119) yields
${\mathbf{E}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\varepsilon_{1}}{{G_{1}}}+\frac{{{{{{G_{1}^{2}}}}}\tilde{\varepsilon}_{2}^{2}{T^{1-c}}}}{{({1-c})}}+\frac{{2{{{p^{2}{G_{1}^{2}}}}}{T^{1-c}}}}{{({1-c})}}+\frac{{R{{(\mathbb{X})}^{2}}{T^{1-c}}}}{{4({1-c})}}+\frac{{r{(\mathbb{X})^{2}}{T^{1-c}}}}{{4({1-c})}}$
$\;\;+\frac{{{{G_{1}}}R(\mathbb{X}){T^{1-c}}}}{{1-c}}+\frac{{{G_{1}}r(\mathbb{X}){T^{1-c}}}}{{1-c}}+\frac{1}{{2n{\alpha_{0}}}}\sum\limits_{i=1}^{n}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}}]$
$\;\;+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{{\varepsilon_{5}}}{\Big{(}{\frac{1}{{{\alpha_{t}}}}-\frac{1}{{{\alpha_{t-1}}}}-\mu}\Big{)}}}{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{\|{\hat{y}-{x_{i,t}}}\|^{2}}]$
$\leq{\varepsilon_{1}}{{G_{1}}}+\frac{{{{{{G_{1}^{2}}}}}\tilde{\varepsilon}_{2}^{2}{T^{1-c}}}}{{({1-c})}}+\frac{{2{{{p^{2}{G_{1}^{2}}}}}{T^{1-c}}}}{{({1-c})}}+\frac{{R{{(\mathbb{X})}^{2}}{T^{1-c}}}}{{4({1-c})}}+\frac{{r{(\mathbb{X})^{2}}{T^{1-c}}}}{{4({1-c})}}$
$\;\;+\frac{{{{G_{1}}}R(\mathbb{X}){T^{1-c}}}}{{1-c}}+\frac{{{G_{1}}r(\mathbb{X}){T^{1-c}}}}{{1-c}}+4{\varepsilon_{5}}{[{1-\mu}]_{+}}{R{{(\mathbb{X})}^{2}}}.$ (120)
Hence, (22) holds.
(ii) From (93), (99), (105), and (116)–(117), we have (23).
(iii) From (109), (111), and (113), and (116)–(117), we have (24).
E. Proof of Theorem 4
We will show that (26)–(28) in Theorem 4 hold in the following, respectively.
Note that (115)–(117) still hold.
(i) From (25) and (118), we have
$\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\big{(}{{\Delta_{i,t}}({\hat{y}}){-\frac{\mu}{2}{{\|{\hat{y}-{x_{i,t}}}\|}^{2}}}}\big{)}}}\leq 0.$ (121)
From (25), we have
[…]
$\sum\limits_{t=1}^{T}{\frac{{\xi_{t}^{2}}}{{{\alpha_{t}}}}}$ $=\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{1}{\mu}\big{(}{\log(T)+1}\big{)},$ (123)
$\sum\limits_{t=1}^{T}{\frac{{\delta_{t}^{2}}}{{{\alpha_{t}}}}}$ $=r{(\mathbb{X})^{2}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{{r{{(\mathbb{X})}^{2}}}}{\mu}\big{(}{\log(T)+1}\big{)},$ (124)
$\sum\limits_{t=1}^{T}{{\xi_{t}}}$ $=\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{1}{\mu}\big{(}{\log(T)+1}\big{)},$ (125)
$\sum\limits_{t=1}^{T}{{\delta_{t}}}$ $=r(\mathbb{X})\sum\limits_{t=1}^{T}{{\alpha_{t}}}\leq\frac{{r(\mathbb{X})}}{\mu}\big{(}{\log(T)+1}\big{)}.$ (126)
Choosing $y={x^{*}}\in{\mathcal{X}_{T}}$, and combining (115), and (121)–(126) yields
${\mathbf{E}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\varepsilon_{1}}{{G_{1}}}+\frac{{{{{G_{1}^{2}}}}\tilde{\varepsilon}_{2}^{2}\big{(}{\log(T)+1}\big{)}}}{{\mu}}+\frac{{2{{p^{2}{G_{1}^{2}}}}\big{(}{\log(T)+1}\big{)}}}{{\mu}}$
$\;\;+\frac{{R{{(\mathbb{X})}^{2}}\big{(}{\log(T)+1}\big{)}}}{{4\mu}}+\frac{{r{{(\mathbb{X})}^{2}}\big{(}{\log(T)+1}\big{)}}}{{4\mu}}$
$\;\;+\frac{{{G_{1}}R(\mathbb{X})\big{(}{\log(T)+1}\big{)}}}{\mu}+\frac{{{G_{1}}r(\mathbb{X})\big{(}{\log(T)+1}\big{)}}}{\mu}.$ (127)
Hence, (26) holds.
(ii) From (25), we have
$\sum\limits_{t=1}^{T}{\frac{1}{{{\gamma_{t}}}}=\sum\limits_{t=1}^{T}{\frac{{{\alpha_{t}}}}{{{\gamma_{0}}}}}\leq}\frac{1}{{{\gamma_{0}}\mu}}\big{(}{\log(T)+1}\big{)},$ (128)
[…]
$\sum\limits_{t=1}^{T}{\frac{1}{{\gamma_{t}^{2}}}}=\sum\limits_{t=1}^{T}{\frac{{\alpha_{t}^{2}}}{{\gamma_{0}^{2}}}}\leq\frac{2}{{\gamma_{0}^{2}{\mu^{2}}}},$ (130)
$\sum\limits_{t=1}^{T}{\xi_{t}^{2}}=\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{2}{{{\mu^{2}}}},$ (131)
$\sum\limits_{t=1}^{T}{\delta_{t}^{2}}=r{(\mathbb{X})^{2}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{2r{{(\mathbb{X})}^{2}}}}{{{\mu^{2}}}},$ (132)
$\sum\limits_{t=1}^{T}{\frac{{{\xi_{t}}}}{{{\gamma_{t}}}}}=\frac{1}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{2}{{{\gamma_{0}}{\mu^{2}}}},$ (133)
$\sum\limits_{t=1}^{T}{\frac{{{\delta_{t}}}}{{{\gamma_{t}}}}}=\frac{{r(\mathbb{X})}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\alpha_{t}^{2}}\leq\frac{{2r(\mathbb{X})}}{{{\gamma_{0}}{\mu^{2}}}},$ (134)
$\sum\limits_{i=1}^{n}\frac{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}}]}{{{2\gamma_{0}}}}+{{\hat{h}}_{T}}({\hat{y}})\leq\sum\limits_{i=1}^{n}\frac{{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}}]}{{{2\gamma_{0}}}}-\sum\limits_{i=1}^{n}\frac{\mu{\mathbf{E}_{{\mathfrak{U}_{t}}}}[{{{\|{\hat{y}-{x_{i,1}}}\|}^{2}}}]}{{{2\gamma_{1}}}}=0.$ (135)
From (117), and (128)–(135), we have
$\;\;\;\;\;{\mathbf{E}}\Big{[}{\Big{(}{\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}}}\Big{)}^{2}}\Big{]}$
$\leq 2G_{2}^{2}{{\tilde{\varepsilon}}_{3}}T+\frac{{2nF{\varepsilon_{2}}T\big{(}{\log(T)+1}\big{)}}}{{{\gamma_{0}}\mu}}+\frac{{4n{{p^{2}{G_{1}^{2}}}}{\varepsilon_{2}}T}}{{{\gamma_{0}}{\mu^{2}}}}+\frac{{nR{{(\mathbb{X})}^{2}}{\varepsilon_{2}}T}}{{2{\gamma_{0}}{\mu^{2}}}}+\frac{{nr{{(\mathbb{X})}^{2}}{\varepsilon_{2}}T}}{{2{\gamma_{0}}{\mu^{2}}}}$
$\;\;+\frac{{2n{G_{1}}R(\mathbb{X}){\varepsilon_{2}}T}}{{{\gamma_{0}}{\mu^{2}}}}+\frac{{2n{G_{1}}r(\mathbb{X}){\varepsilon_{2}}T}}{{{\gamma_{0}}{\mu^{2}}}}.$ (136)
It follows from (136) that (27) holds.
(iii) Choosing $y={x_{s}}$ in (116), and using (15a) and (106) gives
${\varsigma_{s}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}}\leq{{\tilde{h}}_{T}}({\hat{y}})+{{\hat{h}}_{T}}({\hat{y}}).$ (137)
Combining (128)–(135), and (137) gives
$\;\;\;\;\;{\mathbf{E}}\Big{[}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}}\Big{]}$
$\leq\frac{{2nF\big{(}{\log(T)+1}\big{)}}}{{{\gamma_{0}}\mu{\varsigma_{s}}}}+\frac{{4n{{p^{2}{G_{1}^{2}}}}}}{{{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}+\frac{{nR{{(\mathbb{X})}^{2}}}}{{2{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}+\frac{{nr{{(\mathbb{X})}^{2}}}}{{2{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}+\frac{{2n{G_{1}}R(\mathbb{X})}}{{{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}+\frac{{2n{G_{1}}r(\mathbb{X})}}{{{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}.$ (138)
From (25), (67), (137), and (138), we have
$\;\;\;\;\;{\mathbf{E}}[{\rm{Net}\mbox{-}\rm{CCV}}(T)]$
$\leq n{G_{2}}{\tilde{\varepsilon}_{1}}+\frac{{{\varepsilon_{3}}\big{(}{\log(T)+1}\big{)}}}{\mu}+\frac{{2nF{\varepsilon_{4}}\big{(}{\log(T)+1}\big{)}}}{{{\gamma_{0}}\mu{\varsigma_{s}}}}+\frac{{4n{{p^{2}{G_{1}^{2}}}}{\varepsilon_{4}}}}{{{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}+\frac{{nR{{(\mathbb{X})}^{2}}{\varepsilon_{4}}}}{{2{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}$
$\;\;+\frac{{nr{{(\mathbb{X})}^{2}}{\varepsilon_{4}}}}{{2{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}+\frac{{2n{G_{1}}R(\mathbb{X}){\varepsilon_{4}}}}{{{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}+\frac{{2n{G_{1}}r(\mathbb{X}){\varepsilon_{4}}}}{{{\gamma_{0}}{\mu^{2}}{\varsigma_{s}}}}.$ (139)
It follows from (139) that (28) holds.